### v1

# Efficiently Training 7B LLM with 1 Million Sequence Length on 8 GPUs

Abstract.
Nowadays, Large Language Models (LLMs) have been trained using extended context lengths to foster more creative applications. However, long context training poses great challenges considering the constraint of GPU memory. It not only leads to substantial activation memory consumption during training, but also incurs considerable memory fragmentation. To facilitate long context training, existing frameworks have adopted strategies such as recomputation and various forms of parallelisms. Nevertheless, these techniques rely on redundant computation or extensive communication, resulting in low Model FLOPS Utilization (MFU). In this paper, we propose Memo, a novel LLM training framework designed for fine-grained activation memory management. Given the quadratic scaling of computation and linear scaling of memory with sequence lengths when using FlashAttention, we offload memory-consuming activations to CPU memory after each layer’s forward pass and fetch them during the backward pass. To maximize the swapping of activations without hindering computation, and to avoid exhausting limited CPU memory, we implement a token-wise activation recomputation and swapping mechanism. Furthermore, we tackle the memory fragmentation issue by employing a bi-level Mixed Integer Programming (MIP) approach, optimizing the reuse of memory across transformer layers. Empirical results demonstrate that Memo achieves an average of $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively. This improvement is attributed to Memo’s ability to minimize memory fragmentation, reduce recomputation and intensive communication, and circumvent the delays associated with the memory reorganization process due to fragmentation. By leveraging fine-grained activation memory management, Memo facilitates efficient training of 7B LLM with 1 million sequence length on just 8 A800 GPUs, achieving an MFU of 52.30%.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Since the advent of ChatGPT (OpenAI, 2023), Large Language Models (LLMs) have demonstrated remarkable proficiency in comprehending and generating natural language texts. Besides revolutionizing the field of language processing, which encompasses translation (Zhu et al., 2023), coding (Rozière et al., 2023; Guo et al., 2024b), etc., transformer-based LLMs have also found applications in multi-modal scenarios, such as image processing (Dosovitskiy et al., 2021; Peebles and Xie, 2023), video stream analysis (Ruan and Jin, 2022), and AI for science (Bi et al., 2023; DBL, 2023). To accommodate novel applications that require lengthy contexts, LLMs have developed to support long context input, from 2K-4K (Touvron et al., 2023; Taori et al., 2023) to 32K (Jiang et al., 2023; Together.ai, 2023), 128K (OpenAI, 2023; Fu et al., 2024), or even millions of tokens (AI, 2024; Cloud, 2024; Liu et al., 2024). Considering the extrapolation problem (Press et al., 2022; Liu et al., 2023a), which refers to the decline in LLM performance when input sequences exceed the training length, it is necessary to conduct long context training (Jacobs et al., 2023; Chen et al., 2024; Fang and Zhao, 2024) or fine-tuning (Peng et al., 2023; Ding et al., 2024) to facilitate long sequence inference. Beyond natural language processing, increasing the context length is also essential across diverse domains, including video processing (Zheng et al., 2024), protein properties prediction (Chandra et al., 2023), weather forecasting (Nguyen et al., 2023), and health care (Li et al., 2022).
However, training LLMs with long sequence lengths poses a significant challenge for GPU memory. During training, a large amount of activations (i.e., the intermediate results computed in the forward pass) must be stored for gradient computation during the backward pass, resulting in substantial memory consumption. Typically, it is well known that the self-attention module in the transformer architecture has a quadratic computation and memory complexity w.r.t. the sequence length. FlashAttention (Dao et al., 2022; Dao, 2023), now a standard technique for attention computation in LLM training, accelerates computation and shrinks the memory complexity to be linear w.r.t. the sequence length by scheduling memory I/O and recomputing necessary components during the backward pass. Except for attention, the remaining activation memory also scales proportionally with the sequence length, which can become quite large in long context scenarios. For instance, training a GPT model with 7B parameters on a sequence length of 1 million can lead to an activation memory of 4096GB, far exceeding the memory capacity of commonly used accelerators (e.g. 80GB for an NVIDIA H100/A100 GPU).
Moreover, the memory fragmentation issue makes this situation even worse. Besides storing the skeletal activations for the backward pass, there are also tremendous transient activations that are temporarily generated during computation (we will formally categorize the two kinds of activations in Section 3). Such transient activations have distinct data life cycles and usually lead to frequent allocation and deallocation of GPU memory. Currently, most LLM training systems are built on top of PyTorch (Paszke et al., 2019), including Megatron-LM (Shoeybi et al., 2019) and DeepSpeed (Rasley et al., 2020). PyTorch employs a caching memory allocator designed to reduce the costly “cudaMalloc” and “cudaFree” operations by caching and reusing allocated memory blocks. However, the frequent memory (de)allocation requests in the caching allocator result in significant memory fragmentation (Guo et al., 2024a). This issue becomes more severe in long context training, considering the fact that the (de)allocated memory blocks are significantly larger than those in normal tasks. Memory fragmentation not only leads to Out-Of-Memory (OOM) error but also significantly hinders training efficiency because of the frequent triggering of the PyTorch memory reorganization process, which involves calls to expensive “cudaFree” and “cudaMalloc” to release cached blocks and reclaim GPU memory. Figure 1(a) illustrates an example of GPU memory fragmentation. At the peaks of the curves, there is more than 4GB memory reserved but not allocated. However, when the training task tries to allocates 4GB memory, the allocator fails to find a continuous memory space to fulfill the allocation request. Consequently, it necessitates to invoke a series of “cudaFree” and “cudaMalloc” to reorganize memory, which blocks GPU computation.
(a) GPU memory consumption.
(b) One-layer forward time consumption.
Figure 1. The left figure, generated using PyTorch’s snapshot API (Contributors, 2023), shows the allocated and reserved GPU memory of PyTorch when training a 7B GPT model with sequence length 512K. The right figure shows the time consumption of FlashAttention computation, one transformer layer forward computation, and one-layer full activation offloading when training a 7B GPT on 8 A800 GPUs with a TP size of 8.
In this paper, we aim to tackle the memory challenges encountered during long context LLM training. Specifically, we propose and implement an LLM training framework Memo to address the activation data management problem. There are several key observations that inspire our design.
Observation 1: Opportunity for activation swapping. In the field of deep learning training, to reduce the peak memory consumption caused by skeletal activations, activation recomputation  (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) and swapping (Rhu et al., 2016; NVIDIA, 2024b) are two well-known memory reduction techniques that trade time to save memory. [^1] Parallelism techniques like sequence parallelism (Korthikanti et al., 2022; Jacobs et al., 2023) and context parallelism (Liu et al., 2023b; Li et al., 2023; NVIDIA, 2024a) are also compelling approaches to reduce memory at the price of extra communication overhead. Our work is compatible with these parallelism techniques. Typically, both of them reduce memory consumption at the price of extra time cost. For one thing, the activation recomputation technique discards the skeletal activations in the forward pass and later recomputes them in the backward pass, leading to extra computation cost. For another, the swapping technique offloads the activations to CPU memory in the forward pass to relieve the GPU memory pressure, and later fetches them back to GPU memory in the backward pass, incurring the overhead of data transmission between CPU and GPU memory. Contemporary, mainstream LLM training frameworks such as Megatron-LM and DeepSpeed prefer activation recomputation to swapping, which is due to the fact that the GPU computing ability has a far more rapid growth than the connectivity between CPU and GPU memory in the past few years (see Section 2.2 for details).
However, we find that the situation is a bit different in long context training of LLMs. Denote $s$ as the sequence length. The computation complexity of one transformer layer is $O(s^{2})$, while the activation memory complexity is $O(s)$ thanks to FlashAttention. During GPU computation, we can leverage the idle CPU-GPU bandwidth, offloading activations to CPU memory during the forward pass, and fetching the activations during the backward pass. As the sequence length increases, there is greater potential for overlapping computation and communication, given that their time requirements scale quadratically and linearly with the sequence length, respectively. As shown in Figure 1(b), eventually, after reaching a specific sequence length (192K in this case), the transmission of activations can be fully overlapped with GPU computation.
However, in practice, there is limited chance to completely swap all activations. On the one hand, extremely long training data is rare, and most of the time we need to train on data that doesn’t fully overlap the activation transmission and the computation. On the other hand, offloading all activations may cause CPU OOM issues — the CPU memory is responsible for storing all activations from all GPUs on the same machine, but the current CPU memory is typically several terabytes, which is insufficient for very long sequence lengths. Considering the above challenges, we introduce a fine-grained activation recomputation and swapping mechanism to manage the skeletal activations. We consider both tensor-level and token-level activation management. For each layer, following previous works (NVIDIA, 2024b; Li et al., 2023), we consistently offload two kinds of activation tensors, the input of each transformer layer and the output of FlashAttention, to CPU memory. For other activation tensors, we only offload a fraction (denoted as $\alpha$) of tokens, and recompute the rest part during the backward pass. We model the time cost of activation recomputation and transmission and determine the fraction $\alpha$ through a well-formulated linear programmin problem, which aims to maximize offloading activations without impeding GPU computation or causing CPU OOM issues. During the backward pass, prefetching activations can also overlap with GPU computation, because the backward computation is typically twice as much as the forward computation. With both tensor-level and token-level activation management, we make full use of the idle bandwidth and minimize the recomputation overhead to improve the overall efficiency.
Figure 2. An overview of Memo. We devise a fine-grained recomputation and swapping mechanism to manage the skeletal activations for backward propagation, and leverage a bi-level memory planning to reuse the memory space of transient activations across the transformer layers.
Memo Overview
Observation 2: Deterministic memory (de)allocation pattern across iterations and layers. The memory fragmentation mainly comes from frequent and irregular memory (de)allocation requests. However, we observe that, typical LLM training adheres to a deterministic computation process across iterations and layers. All transformer layers in an LLM are identical, and each training iteration involves the same computation. While the general-purpose caching allocator is designed for dynamic computation routines, training LLMs can be conceptualized as static computation graphs (Ansel et al., 2024), which have identical structures across layers. This provides an opportunity to design static planning for each layer and reuse the allocated memory of each layer, thereby mitigating memory fragmentation.
To enhance memory utilization while minimizing fragmentation, we leverage a hierarchical Mixed Integer Programming (MIP) technique to tackle the memory planning problem. Before training, we profile the memory (de)allocation requests of one training iteration, then use MIP to solve an optimized memory plan for a single transformer layer. Since each layer’s memory request are identical, the entire memory block for one layer can be directly reused for the subsequent identical layer. Considering each transformer layer’s memory block as a single memory allocation request, we further solve another MIP problem that plans memory allocation for the entire LLM training, including the initial embedding layer, all transformer layers, and the final classifier layer. We only need to solve the problem once before the actual training, since all iterations can utilize the same memory plan. The near-optimal memory plan eliminates the fragmentation issue and avoids PyTorch’s time-consuming memory reorganization mechanism.
Putting them together, in response to the activation memory challenge in long context training, we propose Memo, an LLM training framework with fine-grained tensor memory management. We consider the challenge as an activation data management problem. To make full use of the idle CPU-GPU bandwidth during training with different sequence lengths, we introduce a token-wise fine-grained activation recomputation and swapping strategy. We employ a bi-level hierarchical MIP technique to solve the memory planning problem and eliminate memory fragmentation. To the best of our knowledge, this is the first training framework that enables efficient training of a 7B LLM on 8 GPUs with a sequence length of 1 million.
Table 1. Commonly used notations in this work.
Notation Explanation
$b$ Batch size
$s$ Context length
$n$ Number of transformer layers
$h$ Hidden size
$P$ Number of model parameters
$\alpha$ The fraction of swapping
We summarize our contributions as follows:
  • We propose and implement an LLM training framework Memo to address the activation data management problem in long context LLM training.
  • We introduce a fine-grained activation recomputation and swapping mechanism to fully utilize the idle CPU-GPU communication bandwidth during time-consuming GPU computation.
  • We employ a bi-level hierarchical MIP technique to solve the memory planning problem and significantly mitigate memory fragmentation.
  • We evaluate Memo through extensive experiments, and demonstrate an average of $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively. Additionally, Memo is the first framework that enables the efficient training of 7B LLM with 1 million context length on only 8 A800 GPUs.
2. Preliminary
In this section, we present an overview of the architecture and training process of LLMs, along with memory reduction strategies and distributed training techniques. Commonly used notations are listed in Table 1.
Figure 3. The architecture of a typical LLM.
gpt_arch
2.1. Large Language Models
2.1.1. Architecture
As shown in Figure 3, the architecture of a typical LLM comprises an input embedding layer, multiple decoder-only transformer layers, and a final classifier layer. The embedding layer converts input tokens into continuous representations. Each decoder-only transformer layer constitutes a multi-head self-attention module with causal mask, and an Feed-Forward Network (FFN) module containing Fully-Connected (FC) networks. The classifier layer takes the hidden states produced by the transformer layers as input, and generates a probability distribution over the vocabulary.
2.1.2. The Training Process
The training process of LLM involves two phases: the forward pass and the backward pass. During the forward pass, the model processes the input data through its layers, and finally generates predictions. The output tensors of the operators in forward pass are called activation tensors, some of which are stored for backward pass computation according to gradient-based learning.
The backward pass, on the other hand, computes the gradients with regard to the model parameters. These gradients are used to update the model’s parameters. Following the chain rule in gradient computation, the backward pass relies on the activation tensors from the forward pass to compute gradients.
2.1.3. The Challenge of Huge Memory Requirement in Long Context Training
In the original form of self-attention, which is the most critical module in LLMs, input tokens are initially projected into queries, keys, and values. The queries and keys are multiplied and then softmaxed to create an $s\times s$ matrix that represents attention weights, which are subsequently used to compute a weighted sum of the values. Thus, it exhibits a memory complexity of $O(s^{2})$.
As the sequence length increases, storing the entire $s\times s$ tensors for the backward pass becomes infeasible. FlashAttention (Dao et al., 2022; Dao, 2023) has emerged as a solution for efficient attention computation in long context training. It minimizes memory I/O and carries out the attention computation in a streaming fashion in both the forward and backward pass, and thereby successfully get rid of the huge $s\times s$ matrix. With such designs, FlashAttention accelerates the overall computation speed (while maintaining a time complexity of $O(s^{2})$), and eliminates the $O(s^{2})$ memory requirement. Currently, FlashAttention has become the de-facto strategy for self-attention computation. Thus, we assume FlashAttention is employed throughout this work.
Although FlashAttention has reduced the memory complexity of LLM training from $O(s^{2})$ to $O(s)$, the linearly scaling activation memory remains the primary challenge in long context training. For example, as we will elaborate in Section 3, when training a 7B GPT model with 32 layers and a hidden size of 4096, using a single 1 million length sequence, the forward activation tensors required by the backward pass consume 4096GB (when using half-precision numbers), whereas the typical memory capacity of a GPU is much smaller. To cope with this issue, there are two lines of efforts, which are the memory reduction techniques and distributed parallelism strategies. In the rest of this section, we will introduce these two lines respectively. It is worth noting that although our work primarily concentrates on the memory reduction techniques, the proposed Memo framework is compatible with a wide range of parallelism strategies.
2.2. Memory Reduction Techniques
Limited GPU memory has become the bottleneck of long context LLM training. To alleviate the memory pressure, there are two notable memory reduction techniques, namely activation recomputation and swapping.
Activation recomputation (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) (a.k.a. activation checkpointing) selectively stores only the inputs of certain layers rather than storing all intermediate activations. During the backward pass, the required activations are recomputed on-the-fly. While this approach reduces the activation memory footprint required for LLM training, it introduces additional computation, which impacts efficiency. Swapping (Rhu et al., 2016; NVIDIA, 2024b; Ren et al., 2021), also known as CPU offloading, aims to relieve the GPU memory pressure by offloading GPU tensors to CPU memory, and fetch them back to GPU when needed. Although swapping does not consumes GPU computation units, it would still slow down the training if the data transmission cannot be overlapped by GPU computation. In general, both the two memory reduction techniques release the memory of activations in the forward pass, but need to rematerialize them in the backward pass, at the price of extra computation or data transmission overhead, respectively.
[…]
2.3. Distributed Parallelism Strategies
Distributed training is essential for efficiently training LLMs, especially in the senario of long context training. To facilitate the training of large-scale data and model, several distributed parallelism strategies have been proposed.
Data Parallelism (DP) (Dean et al., 2012; Zinkevich et al., 2010; Li et al., 2014) duplicates model parameters and distributes the input data across multiple devices. Each device holds a complete copy of the model and processes its input data independently. After backward propagation, the devices synchronize parameter gradients to ensure consistency across the model copies.
Zero-Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2019) is a series of variants built upon DP, aiming to alleviate memory pressure. In naive DP, model parameters, gradients and optimizer states are replicated among all devices. ZeRO is designed in three stages to reduce these memory requirements respectively. First, ZeRO-1 partitions the optimizer states among all DP workers. Next, ZeRO-2 extends ZeRO-1 by also partitioning gradients, further reducing memory footprint. Finally, ZeRO-3, based on ZeRO-2, partitions model parameters among DP workers, further mitigating memory pressure but introducing additional communication to gather parameters during forward and backward pass.
Tensor Parallelism (TP) (Shoeybi et al., 2019) partitions the self-attention and feed-forward modules of transformer layers across multiple devices along either the column or row dimension. It addresses the problem that LLMs can not fit into the memory of a single device. It involves extra collective communication operations (i.e. AllReduce) to synchronize the intermediate results. Therefore, TP is usually applied within a computing node, where intra-node GPUs are connected via high-bandwidth NVLink.
Pipeline Parallelism (PP) (Huang et al., 2019; Narayanan et al., 2021) is also proposed to address the problem that LLMs cannot be fit onto a single device. Different from TP, PP partitions different model layers into several stages, then distributes the stages to different devices. The input data is processed through these stages in a pipeline fashion. Given the peer-to-peer communication style, the PP stages are often distributed across nodes. However, PP introduces a phenomenon known as “bubble”, which corresponds to GPU idle time. The issue becomes more severe when the number of micro-batches is small.
To facilitate efficient long context training, several novel parallelism strategies have been recently proposed.
Sequence Parallelism (SP) (Korthikanti et al., 2022) is built upon TP to further reduce activation memory overhead. It splits the sequence dimension in the part of the model that does not apply TP. The original AllReduce communication now transitions to AllGather and ReduceScatter communication.
DeepSpeed-Ulysses (Jacobs et al., 2023), built upon ZeRO, is another form of sequence parallelism. During self-attention computation, it splits the head dimension, whereas in other model components, it partitions the sequence dimension. For transitioning between modules, it utilizes AllToAll communications, theoretically reducing communication overhead compared to SP. However, its SP degree is limited by the number of heads in self-attention. To further relieve the memory pressure, DeepSpeed-Ulysses leverages ZeRO to distribute model parameters.
Context Parallelism (CP) (Liu et al., 2023b; Li et al., 2023; NVIDIA, 2024a) proposes sharding the query, key, value matrices within the self-attention module along the sequence dimension across different devices. During attention computation, necessary communications are involved to ensure consistent results. The communications can be overlapped with computation by careful scheduling.
In practice, these parallelism strategies and memory reduction techniques can be integrated and employed simultaneously to facilitate efficient training of LLMs.
Figure 4. An example memory request sequence of a transformer layer’s forward and backward propagation. Tensor 15 and 16 are skeletal tensors, while the others are transient tensors.
An example memory sequence
3. Anatomy and System Desiderata
Given the fact that the memory consumption of activations scales proportionally w.r.t. the sequence length in LLM training, we first provide an in-depth anatomy of the key characteristics of different activations in this section. Based on this analysis, we present the design desiderata that motivate the development of Memo.
3.1. Categorization of Activation Tensors
To be specific, according to their life cycles, we categorize activations generated during the forward propagation into two classes, which are the skeletal activations and the transient activations, where the former is necessary for the backward propagation while the latter is not.
For illustration, in Figure 4, tensors 13, 14, 17, 18, and 19 are produced during the forward propagation of a transformer layer, and are discarded before the completion of this layer’s forward pass. Similarly, tensors 20, 21, 22, 23, and 24 are generated during the backward propagation of this layer, and are discarded after corresponding computation. We term them “transient tensors” because they are both created and discarded within a single layer’s forward or backward pass. Transient tensors usually serve as temporary results. Conversely, tensors 15 and 16 are generated during the forward propagation and are needed for backward propagation, so they are discarded in this layer’s backward pass. We refer to these tensors as “skeletal tensors” because they are produced during the forward pass, and are essential for the gradient calculation during the backward pass.
3.2. Analysis of Skeletal Activations
Figure 5. Illustration of the transformer layer architecture. The sizes of skeletal activations are provided in the brackets.
Time consumption
Figure 6. Forward and backward propagation with rounding buffers for token-wise recomputation/swapping. During forward propagation, the darker part in the rounding buffers is offloaded to CPU, while the lighter part is discarded; during backward propagation, the darker part in the rounding buffers if prefetched from CPU, while the lighter part it recomputed.
Rounding buffers with token-wise recomputation
Figure 5 presents all skeletal tensors generated within a transformer layer’s forward propagation, along with their sizes. We can see that the total size of all skeletal activations in a single transformer layer amounts to $16bsh$. To exemplify, when training the GPT-7B model ($h=4096$, 32 layers) with a sequence length ($s$) of 1 million, if we store the skeletal activations in half-precision floating numbers, it would take 4096 GB for only one sequence ($b=1$), exceeding the memory capacity of even 50 A100/H100 GPUs.
The most important characteristic of skeletal activations is that they are needed by backward computation, so they must reside in GPU memory at least before the backward propagation of the corresponding transformer layer begins. However, there is no doubt that maintaining all skeletal activations for backward propagation is infeasible. To this end, memory-saving techniques like activation recomputation and swapping become necessary for long context training. These techniques first release the skeletal activations of a transformer layer in the forward propagation, and later rematerialize them before the corresponding backward propagation. In essence, both the two techniques trade time for memory — the activation recomputation technique incurs extra computation overhead while the swapping technique necessitates transmitting the activations from CPU memory to GPU memory. For extremely long context lengths, using either technique would take significant time to rematerialize the skeletal activations, causing performance degradation.
As a result, we desiderate a meticulous orchestration of the two memory-saving techniques to manage the skeletal activations, so that we can minimize the extra overhead while accommodating the huge memory requirement in long context training of LLMs. To achieve this goal, we develop a token-wise activation recomputation and swapping mechanism, which will be demonstrated in Section 4.1.
3.3. Analysis of Transient Activations
Transient activations are intermediate results generated and discarded during the forward (or backward) pass of a transformer layer. Actually, there are more transient activations than skeletal activations in a transformer layer. Specifically, we observe that the number of transient activations can exceed 5 times that of skeletal activations. Without careful management, the frequent allocation and deallocation can lead to memory fragmentation and degrades system performance.
Fortunately, the most important characteristic of transient activations is that they are identical cross transformer layers, which provides us with the opportunity to manage and reuse their memory regions to minimize fragmentation. In particular, the memory addresses of a single transformer layer’s transient activation tensors can be reused by all other transformer layer’s corresponding transient activation tensors. However, in practice, memory reuse is not fulfilled because the PyTorch caching allocator lack prior information of the memory request sequence during training iterations. This inspires us to statically plan the memory addresses of each transformer layer’s transient tensors, which will be described in detail in Section 4.2.
4. Memo Design
In this section, we propose Memo for fine-grained activation memory management. Our proposed method leverages fine-grained and structured activation management, akin to concise memos that share vital information. The main challenge of long context training is the large activation size which scales linearly w.r.t. sequence length. And we propose token-wise activation recomputation and swapping, along with a bi-level memory planning to address the issue, which targets skeletal activation tensors and transient activation tensors, respectively. The overview of Memo is depicted in Figure 2.
4.1. Token-wise Recomputation and Swapping
Skeletal tensors, generated during the forward pass of a transformer layer, must reside in GPU memory for the subsequent backward propagation. In practice, as sequence length grows, the size of skeletal activations increases linearly, which can easily exceed the capacity of GPU memory. As introduced in Section 2.2, currently the most widely-used technique to tackle this issue is activation recomputation, which stores only the input of each transformer layer, and discard the rest skeletal activation tensors of this layer. Prior to backward propagation of each layer, an additional forward pass of the layer is conducted to reconstruct all skeletal tensors so that the backward computation can be carried out. However, we note that the vanilla activation recomputation strategy is not an optimal choice to handle the challenge of linearly increasing skeletal activation memory, considering the following two reasons: (1) activation recomputation introduces redundant computation, thus diminishing training efficiency; and (2) the memory overhead of retaining the input tensor of each transformer layer can still be expensive, especially when the sequence length is too long or the number of layers is too large. Take the training of GPT-7B with a context length of 1 million as an example again. For only one sequence, the input tensors of all 32 transformer layers together consume 128GB. Even using a SP degree of 8, it takes 16GB for each GPU to store the input tensors of all 32 transformer layers, which already takes up to 20% of total GPU memory capacity.
As explained in Observation 1, the computation complexity of FlashAttention w.r.t. sequence length is $O(s^{2})$, while the size of skeletal activations within a transformer layer scales linearly with sequence length. This provides us with the opportunity to offload skeletal activations to CPU memory, thereby saving GPU memory. We can prefetch them back to GPU before the backward propagation of the corresponding transformer layer. The swapping of skeletal activations can overlap with GPU computations in long context training, since the CPU-GPU data transmission does not consume GPU computation units.
To facilitate the overlapping, we utilize two rounding GPU buffers to store the skeletal activations for all transformer layers. The two rounding buffers are allocated before the actual training iterations begin. As shown in Figure 6, transformer layers with even layer indices place their skeletal activation tensors in rounding buffer 0, while layers with odd layer indices use rounding buffer 1.
After the computation of transformer layer $i$, rounding buffer $(i\%2)$ will be offloaded to CPU using a separate CUDA stream. This happens simultaneously with the computation of transformer layer $(i+1)$. Before the forward computation of transformer layer $(i+2)$, a CUDA event is employed to ensure the content of rounding buffer $(i\%2)$ has been fully offloaded to CPU memory, thus the transformer layer $(i+2)$ can safely rewrite rounding buffer $(i\%2)$.
For backward propagation, after the backward pass of transformer layer $(i+2)$ ends, the contents within rounding buffer $(i\%2)$ become useless, and we start prefetching the skeletal activations of transformer layer $i$ to rounding buffer $(i\%2)$ using another CUDA stream. The prefetching of transformer layer $i$’s skeletal activations happens simultaneously with the backward propagation of transformer layer $(i+1)$. When the sequence length is sufficiently long, with careful computation-transmission overlapping and synchronization, CPU swapping can substitute activation recomputation without incurring additional computation overhead.
However, there are two constraints that prevent us from offloading all skeletal activations to CPU memory.
  • For sequence lengths that are not sufficiently long, the time required to offload all skeletal activations to CPU memory surpasses the computation time for a single transformer layer. This discrepancy forces the computation of transformer layer $(i+2)$ to be delayed until the offloading of rounding buffer to CPU memory is completed, thereby blocks the normal GPU computation workflow. For instance, as illustrated in Figure 1(b), when training a 7B GPT model on 8 GPUs with a TP size of 8, ideal overlap between the computation of a transformer layer and the offloading of its skeletal activations occurs only for sequence lengths exceeding 192K. In practice, the sequence lengths of most training datasets for LLMs are moderate and may be not sufficient to ensure an ideal overlap between computation and transmission.
  • In theory, a longer sequence length provides more opportunities for overlapping CPU offloading with GPU computation. However, in practice, the CPU memory in GPU servers is often limited. For a typical GPU server which has several terabytes CPU memory (e.g. 2TB in our environment), this is insufficient to store all skeletal activations when the sequence length is excessively long or the number of transformer layers is too large. For instance, when training the 7B model on a server equipped with 8 GPUs using a sequence length of 1 million, the skeletal activations amount to a total size of 4096GB, which is double the capacity of CPU memory.
Figure 7. Forward time of FlashAttention and other parts of a transformer layer when training a 7B GPT on 8 GPUs with a TP size of 8.
Forward time of FlashAttention and other parts of a transformer layer
Therefore, instead of simply offloading all skeletal activations to CPU memory, we employ selective activation swapping to ensure perfect overlap of computation and transmission for short sequences as well as to avoid depleting CPU memory for extremely long context lengths. Memo manages to determine the selection of swapping at both the tensor and token granularities, as depicted in Figure 6.
At the tensor granularity, we consider the benefits of leveraging the swapping technique rather than the recomputation technique of different modules. As depicted in Figure 7, FlashAttention constitutes the most substantial portion of the forward computation of a transformer layer. Notably, when the sequence length exceeds 576K, FlashAttention accounts for more than 90% of the computation involved in a single transformer layer. However, as illustrated in Figure 5, the output of FlashAttention only accounts for 6.25% of total skeletal activation size. This inspires us to offload the entire output tensor of FlashAttention to CPU memory since recomputing its output is very time-consuming. Besides, since LLMs have a layered structure, in order to reconstruct the “input_norm”, “q”, “k”, “v” tensors, we also store the input of each transformer layer to CPU, following common recomputation strategy (Chen et al., 2016).
At the token granularity, we develop the token-wise activation recomputation and swapping technique to reduce the memory consumption of all skeletal activation tensors other than the output of FlashAttention and the input of each transformer layer. To be specific, as shown in Figure 6, for each of these skeletal activation tensors, we only offload a fraction (denoted as $\alpha$) to CPU, while the remaining part is discarded, ensuring perfect overlapping and to avoid CPU OOM error. Before the backward propagation, the discarded part is rematerialized via recomputation while the offloaded part is prefetched.
To determine the fraction $\alpha$, we solve the following problem:
(1) $\max\quad$ $\alpha,$
(2) s.t. $(S_{input}+S_{attn}+\alpha\cdot S_{others})/B\leq T_{layer},$
(3) $(n-2)(S_{input}+S_{attn}+\alpha\cdot S_{others})\leq M_{CPU}.$
where $S_{input}$, $S_{attn}$, and $S_{other}$ stand for the size of input tensor, the size of FlashAttention output tensor, the total size of other skeletal activation tensors, respectively, $B$ is the PCIe bandwidth between GPU and CPU, $T_{layer}$ is the forward time of a single transformer layer, $n$ is the total number of transformer layers, and $M_{CPU}$ stands for the capacity of CPU memory. It is worth noting that, the last two transformer layers can initiate the backward pass immediately after the forward pass, obviating the need for swapping. These variables can be easily obtained through profiling before training, so we can determine an appropriate $\alpha$ without much effort.
As a special case, when the determined fraction is 0, skeletal buffers except layer input and attention output can share one GPU buffer, as they are fully recomputed and do not need to be offloaded. Thus there is no need to use two rounding buffers to avoid data corruption.
4.2. Bi-level Memory Planning
Figure 8. Bi-level MIP algorithm.
bi-level
In the previous subsection, we have tackled the management of skeletal activations by the fine-grained recomputation and swapping technique. However, frequent allocation and deallocation of the transient activation tensors still lead to GPU memory fragmentation, which forces the allocator to frequently reorganize GPU memory using time-consuming “cudaFree” and “cudaMalloc” operations. To address the issue, and to achieve full reuse of GPU memory across all transformer layers, we design a bi-level Mixed Integer Programming (MIP) method.
In practice, our initial step involves profiling the sequence of memory requests during a single training iteration. Given the memory request sequence, the challenge lies in determining the address of each requested tensor while at the same time minimizing the peak memory usage. This task aligns with the well-established offline Dynamic Storage Allocation (DSA) problem (Sekiyama et al., 2018), which can be formulated as a Mixed Integer Programming (MIP) problem. A concise overview of this formulation is shown as follows.
The offline DSA problem assumes a predefined sequence of memory allocations and deallocations, and aims to determine the address of each allocated memory block and at the same time minimizing the peak memory usage. Parameters of offline DSA problem includes:
  • $n$, the number of requested tensors.
  • $S_{i}$, the size of requested tensor $i$, for $\forall i\in\{1,2,...,n\}$.
  • $E=\{(i,j)|\text{tensor }i,j\text{ have overlapped lifespan}\}$.
And the problem can be written as
[…]
where $A_{i}$ stands for the address of requested tensor $i$, $M$ stands for the peak memory usage, $M_{cap}$ is the memory capacity, and $z_{ij}$ is defined as
[…]
Here the first constraint and the last two constraints define and limit peak memory, while the second and third constraints ensure non-overlapping tensors. Following this formulation, the solution for each tensor’s address is optimal. However, modern LLM training involves thousands of allocation and deallocation requests, which makes this MIP problem computationally intractable.
Fortunately, all transformer layers have identical structures and memory request sequences, which presents repetitive substructures within the MIP problem. By leveraging this inherent repetitiveness, we can devise a bi-level hierarchical MIP optimization algorithm, which is both computationally feasible and effective.
Figure 9. Memory request sequence during training.
transient request sequence
As discussed in Section 2.1.1, a typical LLM consists of an embedding layer, $n$ consecutive transformer layers, and a final classification layer. As shown in Figure 9, each layer has forward memory request sequence and backward memory request sequence. The memory request sequence is in the form of a sequence of “malloc tensor_id size” and “free tensor_id size”. Since all transformer layers in an LLM are identical, they have the same forward/backward pass memory request sequence.
As shown in the bottom of Figure 8, we first solve the offline DSA sub-problem for just one transformer layer’s forward (backward) pass, which is called the first-level MIP. This offline DSA problem can be simply solved by any MIP solver. After this step, the peak memory needed for the forward (backward) propagation of a single transformer layer, as well as the address of each transient tensor within a transformer layer is determined. After solving the sub-problem for one transformer layer, all other transformer layers can reuse the same memory address for (de)allocation.
Subsequently, we can replace the original fine-grained memory request sequence of a transformer layer forward (backward) propagation with a “pseudo” large memory request pair, as shown in Figure 8. After the substitution, this reformulated memory request sequence also satisfies the formulation of an offline DSA problem, with a size small enough to be efficiently solved. We then leverage the MIP solver again to solve this second-level MIP problem. After this step, the addresses of all activation tensors, as well as the peak memory needed for all transient activation tensors can be determined.
4.3. System Implementation
Figure 10. Overall architecture of Memo.
Overall architecture of our system
4.3.1. Overview
Figure 10 illustrates the overall architecture of Memo. First, the job profiler takes in the model configuration and user-defined settings, then executes a training iteration to profile the memory requests directed to the PyTorch CUDA allocator during the training phase. The job profiler also determines offloading fraction $\alpha$ by solving optimization problem in Section 4.1. These memory requests comprise a sequence of allocation and deallocation instructions. Afterwards, the memory planner receives the memory requests, executes the bi-level MIP optimization algorithm and, generates a memory plan, which constitutes the addresses of all activation tensors during one training iteration. Finally, the runtime executor reads the memory plan and conducts the training process.
4.3.2. Job Profiler
The job profiler is designed to profile the memory request sequence during a training iteration. To implement the module, we have extended the PyTorch CUDA allocator with extra interfaces that log each memory request it receives, in the format of “malloc tenosr_id size” and “free tensor_id size”.
However, naively recording all memory requests may lead to OOM error. For example, profiling a GPT-7B model with a sequence length of 512K on 8 GPUs can result in OOM error since the proposed techniques are not employed in profiling. Fortunately, as the transformer layers have identical memory footprint, we leverage this property by only profiling a single transformer layer’s memory footprint and then applying it to all transformer layers.
When the sequence is too long, we cannot even profile one single transformer layer. In such extreme cases, we turn to the CUDA Unified Memory feature, which enables the swapping between GPU memory and CPU memory under the hook, effectively creating an illusion of unlimited GPU memory. By integrating CUDA Unified Memory support into the PyTorch CUDA allocator, we have successfully managed to profile the training of extremely long context lengths without encountering OOM errors.
The profiler also gathers the basic information to determine $\alpha$ in Section 4.1, including the size of each skeletal activation tensor, and the forward time of a single transformer layer. Subsequently, it solves for the optimal $\alpha$ to maximize the overlapping of computation and tranmission as well as to avoid CPU OOM error.
4.3.3. Memory Planner
Given the memory request sequence generated by the job profiler, memory planner executes the bi-level MIP optimization algorithm as introduced in Section 4.2 to generate a memory plan, which includes the address of each activation tensor and the peak memory usage needed during training. In all our experiments, memory planning takes less than 5 minutes, which is negligible compared to the training time of LLMs.
Figure 11. Scheduling of computation, offloading and prefetching w/ and w/o token-wise recomputation. Given the superior computing ability of modern GPUs, the recomputation part is faster than the offloading part that blocks forward computation.
scheduling of computation, offloading and prefetching during training, without token-wise recomputation.
4.3.4. Runtime Executor
The runtime executor takes the memory plan as input, and executes the training process. It is built on the top of Megatron-LM (Shoeybi et al., 2019) and TransformerEngine (NVIDIA, 2024b), one of the most popular LLM training frameworks. The runtime executor utilizes two rounding buffers for the storage of skeletal activations, as introduced in Section 4.1. Meanwhile, the transient activation tensors are allocated and discarded according to the memory plan.
Three CUDA streams are employed for efficient overlapping of data transmission and GPU computation, which are for GPU computation, activation offloading from GPU to CPU, and activation prefetching from CPU to GPU, respectively. Figure 11 shows the scheduling of computation and transmission. After the computation of a transformer layer’s forward pass, the skeletal activations of this layer are scheduled to be transferred to the CPU memory, which can overlap with the computation of the next layer. Before the backward computation of a transformer layer, the forward skeletal activations of the previous layer are scheduled to be fetched back to GPU. In addition, token-wise tensor recomputation is also scheduled before the layer’s backward pass. By hiding the activation swapping with computation and enabling the fractional, token-wise activation recomputation, Memo minimizes the overhead of activation rematerialization at full stretch.
5. Experiments
In this section, we conduct experiments across various model sizes and input sequence lengths to show that Memo achieves superior efficiency in longer context training of LLMs.
5.1. Setup
Hardware: Our experiments are conducted on an A800 GPU cluster, with each node equipped with 8 NVIDIA A800 GPUs (80GB). The GPUs within each node are interconnected via NVLinks (400GB/s), while the nodes are interconnected through Infiniband (200GB/s). Each node has 2TB CPU memory, and the GPU-CPU communication bandwidth is 32GB/s.
Baselines: We select two widely-used LLM training frameworks as baselines for our experiments. The first is Megatron-LM (commit id: ccfeda47cb) (Shoeybi et al., 2019) in conjunction with TransformerEngine (v1.3) (NVIDIA, 2024b). Megatron-LM, maintained by NVIDIA, is renowned for its comprehensive support of hybrid parallelisms, including DP, TP, PP, SP, and CP. The other baseline is Megatron-Deepspeed (commit id 7eb36a11b3) paired with DeepSpeed (v0.14.3) (Rasley et al., 2020), which is recognized for ZeRO optimizers and DeepSpeed-Ulysses (Jacobs et al., 2023), a novel parallel training strategy designed for long context LLM training.
Metrics: We use two important evaluation metrics to measure the training efficiency, which are Model FLOPs Utilization (MFU) and Tokens per GPU per Second (TGS). MFU is defined as the ratio of model FLOPs per second to the theoretical peak FLOPs per second of the GPU (e.g. 312 TFLOPS for NVIDIA A800 GPUs) (Korthikanti et al., 2022). Based on FlashAttention (Dao, 2023) and considering the causal mask, the exact formula for calculating model FLOPS per sample is:
$6\cdot s\cdot P+6\cdot n\cdot h\cdot s^{2}.$
MFU is a standard metric that measures the training efficiency of how model FLOPs utilize computational resources. On the other hand, TGS directly measures training throughput, providing a clear view of how quickly a model can be trained using a given amount of training samples. Both metrics are crucial for LLM researchers and engineers, enabling comparisons among various training strategies (including distributed parallelisms and activation recomputation).
Table 2. Configurations of the evaluated models.
Hyper Parameters
Model Size $n_{layers}$ $h$ $h_{ffn}$ $n_{head}$ $n_{vocab}$
7B 32 4096 16384 32 50257
13B 40 5120 20480 40 50257
30B 48 7168 28672 56 50257
65B 80 8192 32768 64 50257
Workloads: Our experiments cover a wide range of workloads to examine the strength of Memo. In particular, we consider training the 7B, 13B, 30B and 65B GPT models on 8, 16, 32, and 64 GPUs, respectively, with various sequence lengths ranging from 64K to 1408K. The detailed model configurations are shown in Table 2.
Table 3. MFU (Model FLOPS Utilization) and TGS (Throughput per GPU per Second) of different methods on different number of GPUs, model sizes, and sequence lengths. “DS” stands for “DeepSpeed”, and “Mega” stands for “Megatron-LM”. ✗oom stands for Out Of Memory error when the system runs out of GPU memory, and ✗oohm stands for Out Of Host Memory error, which means the CPU memory is depleted.
Sequence Length
# GPUs Model Size Method 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
27.95% 25.46% 23.38% oom oom oom oom oom oom oom oom oom
DS 935.76 555.51 296.48 oom oom oom oom oom oom oom oom oom
41.55% 24.13% 29.07% 27.98% 34.43% 30.90% oom oom oom oom oom oom
8 7B Mega 1417.74 526.62 368.65 250.07 237.56 173.63 oom oom oom oom oom oom
52.34% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71% 52.30% oom oom oom
Memo 1786.22 1111.99 679.92 474.02 357.70 295.51 245.76 215.98 188.73 oom oom oom
27.97% 25.45% 21.98% oom oom oom oom oom oom oom oom oom
DS 553.85 333.51 171.82 oom oom oom oom oom oom oom oom oom
38.51% 23.02% 25.30% 22.88% 29.10% 19.41% oom oom oom oom oom oom
16 13B Mega 762.46 301.67 197.78 127.40 125.86 83.94 oom oom oom oom oom oom
52.65% 50.93% 51.22% 51.91% 52.40% 52.13% 51.71% 51.76% 52.06% 51.74% 51.78% 52.10%
Memo 1042.5 667.41 400.39 289.11 226.65 184.33 154.63 134.08 118.96 105.74 95.72 87.93
29.93% 25.54% oom oom oom oom oom oom oom oom oom oom
DS 296.41 176.92 oom oom oom oom oom oom oom oom oom oom
35.76% 14.70% 17.15% 23.32% oom oom oom oom oom oom oom oom
32 30B Mega 354.20 101.85 74.24 73.37 oom oom oom oom oom oom oom oom
52.12% 49.66% 50.00% 50.69% 51.06% 51.72% 51.18% 51.50% 51.24% 51.73% 51.59% oohm
Memo 516.16 344.09 216.41 159.52 126.23 105.28 88.55 77.48 68.20 61.72 55.79 oohm
31.05% 26.13% 22.07% 20.40% 19.83% 19.06% 19.53% 19.12% 19.00% 19.11% 18.90% oom
DS 149.8 90.15 48.51 32.89 25.24 20.04 17.50 14.93 13.14 11.86 10.64 oom
22.79% 15.10% 9.57% 12.07% 5.32% oom oom oom oom oom oom oom
64 65B Mega 109.94 52.10 21.03 19.46 6.77 oom oom oom oom oom oom oom
47.80% 48.61% 49.87% 48.85% 49.71% 50.05% 51.16% 51.05% 51.27% 51.20% 51.42% 51.45%
Memo 230.62 167.69 109.58 78.76 63.29 52.65 45.84 39.85 35.45 31.77 28.94 26.50
5.2. End-to-end Evaluation
We compare the end-to-end training efficiency of Memo and two baselines. Table 3 shows the MFU and TGS of DeepSpeed-Ulysses, Megatron-LM and Memo under different training workloads. During evaluation, we manually adjust the distributed parallelism strategies for each system and each workload to achieve optimal training performance for fair comparisons. Detailed parallelism strategies are provided in Appendix A.
Overall, Memo is capable of training longer sequences than the competitors. Across the training of 7B, 13B, 30B, and 65B models on 8, 16, 32, and 64 GPUs, DeepSpeed-Ulysses supports sequence lengths of 256K, 256K, 128K, and 1280K, while Megatron-LM supports sequence lengths of 640K, 640K, 384K, 512K. In comparison, Memo achieves superior performance in all scenarios, enabling training sequence lengths of 1024K, 1408K, 1280K, and 1408K. Megatron-LM only supports up to 640K sequence length, even if we have leveraged a high model parallel degree and enabled the memory reduction techniques. This is unsurprising since it overlooks the memory fragmentation issue, leading to OOM for large sequence lengths. DeepSpeed, thanks to its support of DeepSpeed-Ulysses sequence parallel and ZeRO-3 optimizer, is capable of training 1280K sequence length when training the 65B model on 64 GPUs. When training smaller models, DeepSpeed supports only very small sequence lengths. This is because it can only utilize a small SP size of 8, which either aligns to the number of GPUs or is dividable by the number of attention heads (40 and 56). In contrast, by token-wise recomputation/swapping and memory planning, Memo is able to train sequences over 1 million tokens in all scenarios.
Furthermore, when comparing Memo to the baselines with aligned sequence lengths, Memo achieves superior MFU and TGS. Across all experimented workloads, Memo achieves an average MFU of 51.33%. In contrast, Megatron-LM and DeepSpeed only achieve an average MFU of 23.91% and 23.26%, respectively. In average, Memo achieves $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
The deficiencies of baselines are not surprising — due to their unsatisfactory memory management, Megatron-LM and DeepSpeed require high model parallel size (i.e. large TP size and/or large SP size) and full activation recomputation to avoid OOM error, which leads to low training efficiency. For example, in order to train the 65B model with 1024K sequence length on 64 GPUs, DeepSpeed has to simultaneously use an SP size of 64, ZeRO-3 and activation recomputation to avoid OOM error. The large SP/TP size and ZeRO degree introduce significant communication overheads, while vanilla memory reduction techniques incurs extra overhead, further diminishing training efficiency. Memory fragmentation is another factor leading to the low training efficiency of these systems. When the GPU memory becomes highly fragmented and cannot allocate sufficient space for new tensors, the allocator will call expensive “cudaMalloc” and “cudaFree” operations to reorganize the GPU memory, which blocks GPU computation substantially. When training the 7B model on 8 GPUs using Megatron-LM, the memory reorganization operation is triggered 6 times and 16 times per iteration for sequence lengths of 128K and 256K, respectively.
In contrast, Memo successfully addresses these issues. On the one hand, the fine-grained activation recomputation and swapping technique significantly reduces computation burden compared to the vanilla full recomputation. This advantage is even more pronounced during long-context training. On the other hand, Memo adopts an optimized memory planning for activation, which avoids GPU memory fragmentation and the time-consuming memory reorganization process. Reducing memory fragmentation also allows for more efficient model parallelism configurations. For example, when training the 30B model on 32 GPUs with 256k sequence length, Memo adopts a TP size of 8, a CP size of 2, and a DP size of 2, which is more efficient than Megatron-LM with a TP size of 8 and CP size of 4. Additionally, Memo does not trigger any memory reorganization since the memory is already managed. As a result, Memo consistently achieves an MFU of approximately 50% across all model sizes and sequence lengths, enabling the efficient training of significantly longer sequences compared to the baselines.
Table 4. MFU of different methods for ablation studies. The experiments are conducted by training the 7B model on 8 GPUs.
Sequence Length
Method 64K 128K 256K 384K 512K 640K 768K 896K
Full Recomputation 41.19% 23.00% 29.07% 25.67% oom oom oom oom
Full Recomputation + Memory Plan 42.91% 43.17% 42.05% 42.49% 41.90% 42.15% oom oom
Full Swapping + Memory Plan 37.40% 46.33% 53.62% oohm oohm oohm oohm oohm
Memo (Fine-grained Recomputation & Swapping + Memory Plan) 47.99% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71%
(a) Longest supported sequence length.
(b) MFU at longest sequence length.
(c) MFU at varying sequence lengths.
(d) Training loss curves.
Figure 12. (a), (b): The longest supported sequence length and corresponding MFU of DeepSpeed, Megatron-LM and Memo when training the 7B model on various numbers of GPUs. (c): The MFU when training the 7B model on 64 GPUs with sequence length varying from 1024K to 8192K. (d): Training loss curves of Megatron-LM and Memo with different $\alpha$.
5.3. Ablation Studies
Next, we assess the effectiveness of the proposed techniques in Memo. All experiments in ablation studies are conducted by training the 7B model on 8 GPUs, keeping the parallelism configuration fixed at a TP size of 4 and a CP size of 2.
5.3.1. Effectiveness of Memory Planning
To evaluate the effectiveness of memory planning, we evaluate two variants of Memo with full recomputation, both with and without memory planning. As shown in the first two rows of Table 4, without memory planning, the longest sequence supported is only 384k, achieving an MFU of 25.67%. After applying memory planning, the longest supported sequence length increases to 640k, with an MFU of 42.15%. The results are reasonable since full recomputation without memory planning has severe memory fragmentation, resulting in OOM errors in large sequence length scenarios.
Additionally, the frequent GPU memory reorganization process further impairs training efficiency. By employing memory planning, the fragmentation issue can be minimized, providing more memory for longer context training. Getting rid of GPU memory reorganization, memory planning brings an average of $1.51\times$ MFU when facing the same context length.
5.3.2. Effectiveness of Token-wise Recomputation
For token-wise recomputation and swapping, we compare Memo and its variants, one with full recomputation and another with full swapping. The results are shown in the last two rows of Table 4. When training with appropriate sequence length, which is 256k in this scenario, the computation time of one transformer layer can fully overlap with the offloading time of a layer’s activations. Therefore, full swapping with memory planning can achieve an MFU of 53.62% under 256K sequence length, far exceeding the 42.05% achieved by full recomputation with memory planning. However, for short sequence lengths, such as 64K, the offloading time of one layer’s activations block the GPU computation, resulting in a lower MFU than full recomputation. Full swapping presents another challenge as the sequence length grows longer: the host memory is rapidly depleted by offloaded activations, leading to OOHM errors.
By employing token-wise recomputation together with swapping, Memo consistently improves training efficiency for both short and long context lengths. For short sequence lengths like 64K, our tensor-level design only offloads the input tensor of the transformer layer and the FlashAttention output tensor to CPU memory, enabling efficient overlap of GPU computation and data transmission. For long context lengths, our token-level management successfully avoids depleting the CPU memory, and incurs only minimal recomputation overhead. Among all the methods, Memo supports the longest sequence length. Considering MFU, Memo achieves an average of $1.22\times$ MFU compared to full recomptutation with memory planning, and an average of $1.13\times$ MFU compared to full offloading with memory planning.
5.4. Scalability
To demonstrate the scalability of Memo, we train the 7B model on 8, 16, 32, and 64 GPUs respectively, and report the maximum supported sequence length. As shown in Figure 12(a), when the number of GPUs increases, the maximum sequence length supported by Memo increases linearly. When training on 8, 16, 32, and 64 GPUs, Memo is capable of training 1, 2, 4, 8 million sequence lengths, respectively, which demonstrates ideal scalability. Memo also consistently maintains an MFU of over 50% across different numbers of GPUs, as shown in Figure 12(b).
For DeepSpeed, as the number of GPUs increases, it can enlarge the SP size, leading to longer supported sequence lengths. Note that since because the 7B model has 32 attention heads and thereby the maximum SP size is 32, DeepSpeed achieves the same maximum sequence length 1536K for both 32 and 64 GPUs. Megatron-LM supports context parallelism, which has better scalability than DeepSpeed. When the number of GPUs increases, the longest sequence length it can handle grows sublinearly. Compared to the baselines, Memo introduces fine-grained activation memory management, achieving not only ideal scalability but also better efficiency.
We also evaluate the MFU metrics of the three systems when training the 7B model on 64 GPUs with sequence lengths varing from 1024K to 8192K. In Figure 12(c), as sequence length increases, the MFU of Memo maintains above 50% consistently, surpassing the competitors significantly.
5.5. Convergence of Memo
To demonstrate the correctness of our system implementation, we conduct a convergence experiment. Specifically, we train the 7B model with 128K sequence length on 8 GPUs and compare the convergence of Memo and Megatron-LM for 1000 iterations. For both systems, we fix the parallelism strategy to TP size 4 and CP size 2, and for Memo, we enumerate the value of $\alpha$ in $\{0,0.125,0.25,0.5,1\}$. As shown in Figure 12(d), the loss curves of Memo with different $\alpha$ values all align with Megatron-LM, confirming the correctness of our system implementation.
6. Related Work
Parallelism strategies for long context training: To tackle the challenge of long context training, DeepSpeed-Ulysses (Jacobs et al., 2023) employs an novel AllToAll communication to facilitate the partition of input sequence among GPUs, achieving lower communication overhead compared with Megatron-LM sequence parallelism (Korthikanti et al., 2022). LightSeq (Li et al., 2023), Ring Attention (Liu et al., 2023b), and Megatron-LM context parallelism (NVIDIA, 2024a) propose to split the sequence within self-attention computation, achieving better scalability. Recent efforts in the realm of sequence and context parallelisms (Chen et al., 2024; Fang and Zhao, 2024) aim to integrate multiple strategies and enhance existing distributed settings. It is worth noting that the fine-grained memory management of Memo is orthogonal to these distributed parallelism strategies.
Activation recomputation and swapping: Capuchin (Peng et al., 2020) proposes to combine recomputation and swapping to reduce memory footprint during training. The swapping decision is made by considering the tensor access pattern. In addition to tensor recomputation, MegTaichi (Hu et al., 2022) also proposes to co-optimize the tensor partition. Coop (Zhang et al., 2023) notices that naive tensor recomputation leads to severe memory fragmentation, and proposes heuristics to reduce memory fragmentation during tensor recomputation. While these works offer solutions for common deep learning models, they do not take advantage of the specific characteristics of LLM training to achieve full overlapping and fragmentation minimization.
Memory planning for deep learning models: The memory allocation problem of in deep learning models can be regarded as a DSA problem and solved by MIP (Sekiyama et al., 2018).OLLA (Steiner et al., 2022) proposes to optimize the lifetime and memory location of tensors during the training process by solving a joint ILP problem, reducing the peak memory during training. However, these methods does not exploit the repetitive substructure in LLMs and relies on heuristics to simplify the integer programming problem.
7. Conclusion
In this paper, we proposed Memo to address the memory challenges in long context LLM training. We designed a fine-grained activation recomputation and swapping strategy to fully utilize the idle PCIe bandwidth during the GPU computation, thereby reducing the activation rematerialization cost in long context LLM training. We employed a bi-level MIP technique to solve the problem of memory allocation within one transformer layer, and reused the same memory space for each identical layer so as to eliminate memory fragmentation. Through extensive experiments, we demonstrated that Memo achieved an average of $2.42\times$ MFU compared to Megatron-LM. By leveraging fine-grained tensor memory management, Memo achieved 52.30% MFU when training 7B LLM with 1 million sequence length on only 8 A800 GPUs.
Appendix A Detailed Parallelism Strategy in Evaluation
Table 56, and 7 present the detailed parallelism training strategies for end-to-end evaluation.
Table 5. Detailed parallelism strategies for DeepSpeed. “SP” stands for DeepSpeed-Ulysses sequence parallel size, “DP” stands for data parallel size, “ZeRO” stands for ZeRO stage, “AR” stands for activation recomputation.
Sequence Length
# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
SP 2 4 8 oom oom oom oom oom oom oom oom oom
DP 4 2 1 oom oom oom oom oom oom oom oom oom
ZeRO 3 3 3 oom oom oom oom oom oom oom oom oom
8 7B AR On On On oom oom oom oom oom oom oom oom oom
SP 4 8 8 oom oom oom oom oom oom oom oom oom
DP 4 2 2 oom oom oom oom oom oom oom oom oom
ZeRO 3 3 3 oom oom oom oom oom oom oom oom oom
16 13B AR On On On oom oom oom oom oom oom oom oom oom
SP 8 8 oom oom oom oom oom oom oom oom oom oom
DP 4 4 oom oom oom oom oom oom oom oom oom oom
ZeRO 3 3 oom oom oom oom oom oom oom oom oom oom
32 30B AR On On oom oom oom oom oom oom oom oom oom oom
SP 8 8 16 32 32 32 64 64 64 64 64 oom
DP 8 8 4 2 2 2 1 1 1 1 1 oom
ZeRO 3 3 3 3 3 3 3 3 3 3 3 oom
64 65B AR On On On On On On On On On On On oom
Table 6. Detailed parallelism strategies for Megatron-LM. “TP” stands for tensor parallel size, “CP” stands for context parallel size, “DP” stands for data parallel size, “PP” stands for pipeline parallel size, “AR” stands for activation recomputation. For all evaluations of Megatron-LM, Megatron-style sequence parallel and ZeRO-1 optimizer are enabled.
Sequence Length
# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
TP 2 4 4 4 oom oom oom oom oom oom oom oom
CP 4 2 2 2 oom oom oom oom oom oom oom oom
DP 1 1 1 1 oom oom oom oom oom oom oom oom
$P$ 1 1 1 1 oom oom oom oom oom oom oom oom
8 7B AR On On On On oom oom oom oom oom oom oom oom
TP 8 4 4 4 8 8 oom oom oom oom oom oom
CP 2 2 2 2 2 2 oom oom oom oom oom oom
DP 1 1 1 1 1 1 oom oom oom oom oom oom
$P$ 1 2 2 2 1 1 oom oom oom oom oom oom
16 13B AR On On On On On On oom oom oom oom oom oom
TP 4 4 8 8 oom oom oom oom oom oom oom oom
CP 4 8 4 4 oom oom oom oom oom oom oom oom
DP 1 1 1 1 oom oom oom oom oom oom oom oom
$P$ 2 1 1 1 oom oom oom oom oom oom oom oom
32 30B AR On On On On oom oom oom oom oom oom oom oom
TP 8 32 16 16 64 oom oom oom oom oom oom oom
CP 8 1 1 4 1 oom oom oom oom oom oom oom
DP 1 2 1 1 1 oom oom oom oom oom oom oom
$P$ 1 1 4 1 1 oom oom oom oom oom oom oom
64 65B AR On On On On On oom oom oom oom oom oom oom
Table 7. Detailed parallelism strategies for Memo. “TP” stands for tensor parallel size, “CP” stands for context parallel size, “DP” stands for data parallel size, “PP” stands for pipeline parallel size, $\alpha$ stands for CPU offload fraction. For all evaluations of Memo, Megatron-style sequence parallel and ZeRO-1 optimizer are enabled.
Sequence Length
# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
TP 2 4 4 4 4 4 4 4 8 oom oom oom
CP 2 2 2 2 2 2 2 2 1 oom oom oom
DP 2 1 1 1 1 1 1 1 1 oom oom oom
$P$ 1 1 1 1 1 1 1 1 1 oom oom oom
8 7B $\alpha$ 0.0 0.0 1.0 0.5 0.125 0.125 0.0 0.0 0.0 oom oom oom
TP 2 4 4 4 4 4 4 4 4 4 4 4
CP 2 2 2 2 4 4 4 4 4 4 4 4
DP 4 2 2 1 1 1 1 1 1 1 1 1
$P$ 1 1 1 2 1 1 1 1 1 1 1 1
16 13B $\alpha$ 0.0 0.0 0.0 0.0 0.5 0.125 0.125 0.0 0.0 0.0 0.0 0.0
TP 4 8 8 8 8 8 8 8 8 8 8 oohm
CP 1 1 2 2 2 4 4 4 4 4 4 oohm
DP 8 4 2 2 2 1 1 1 1 1 1 oohm
$P$ 1 1 1 1 1 1 1 1 1 1 1 oohm
32 30B $\alpha$ 0.0 0.0 0.0 0.0 0.0 0.25 0.125 0.125 0.125 0.125 0.0 oohm
TP 8 4 8 8 8 8 8 8 8 8 8 8
CP 1 4 4 8 8 8 8 8 8 8 8 8
DP 8 4 2 1 1 1 1 1 1 1 1 1
$P$ 1 1 1 1 1 1 1 1 1 1 1 1
64 65B AR 0.0 0.0 0.0 0.25 0.25 0.25 0.25 0.125 0.125 0.0 0.0 0.0

### v2

@@
-Efficiently Training 7B LLM with 1 Million Sequence Length on 8 GPUs
+Memo: Fine-grained Tensor Management For Ultra-long Context LLM Training
 Abstract.
@@
 To maximize the swapping of activations without hindering computation, and to avoid exhausting limited CPU memory, we implement a token-wise activation recomputation and swapping mechanism.
-Furthermore, we tackle the memory fragmentation issue by employing a bi-level Mixed Integer Programming (MIP) approach, optimizing the reuse of memory across transformer layers.
-Empirical results demonstrate that Memo achieves an average of $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
+Furthermore, we tackle the memory fragmentation issue by employing a bi-level Mixed Integer Programming (MIP) approach, optimizing memory reuse across transformer layers.
+Empirical results demonstrate that Memo achieves an average of $1.97\times$ and $1.80\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
 This improvement is attributed to Memo’s ability to minimize memory fragmentation, reduce recomputation and intensive communication, and circumvent the delays associated with the memory reorganization process due to fragmentation.
 By leveraging fine-grained activation memory management, Memo facilitates efficient training of 7B LLM with 1 million sequence length on just 8 A800 GPUs, achieving an MFU of 52.30%.
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+Large Language Model, Long Context Training, Tensor Management, Data Rematerialization
+copyright: acmlicensedjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 1 (SIGMOD)article: 53publicationmonth: 2doi: 10.1145/3709703isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Distributed computing methodologies
 1. Introduction
 Since the advent of ChatGPT (OpenAI, 2023), Large Language Models (LLMs) have demonstrated remarkable proficiency in comprehending and generating natural language texts.
-Besides revolutionizing the field of language processing, which encompasses translation (Zhu et al., 2023), coding (Rozière et al., 2023; Guo et al., 2024b), etc., transformer-based LLMs have also found applications in multi-modal scenarios, such as image processing (Dosovitskiy et al., 2021; Peebles and Xie, 2023), video stream analysis (Ruan and Jin, 2022), and AI for science (Bi et al., 2023; DBL, 2023).
-To accommodate novel applications that require lengthy contexts, LLMs have developed to support long context input, from 2K-4K (Touvron et al., 2023; Taori et al., 2023) to 32K (Jiang et al., 2023; Together.ai, 2023), 128K (OpenAI, 2023; Fu et al., 2024), or even millions of tokens (AI, 2024; Cloud, 2024; Liu et al., 2024).
-Considering the extrapolation problem (Press et al., 2022; Liu et al., 2023a), which refers to the decline in LLM performance when input sequences exceed the training length, it is necessary to conduct long context training (Jacobs et al., 2023; Chen et al., 2024; Fang and Zhao, 2024) or fine-tuning (Peng et al., 2023; Ding et al., 2024) to facilitate long sequence inference.
+Besides revolutionizing the field of language processing, which encompasses translation (Zhu et al., 2024), coding (Rozière et al., 2023; Guo et al., 2024b; Zhang et al., 2024b), etc., transformer-based LLMs have also found applications in multi-modal scenarios, such as image processing (Dosovitskiy et al., 2021; Peebles and Xie, 2023), video stream analysis (Ruan and Jin, 2022), and AI for science (Bi et al., 2023; AI4Science and Quantum, 2023).
+To accommodate novel applications that require lengthy contexts (Zhao et al., 2024), LLMs have developed to support long context input, from 2K-4K (Touvron et al., 2023; Taori et al., 2023) to 32K (Jiang et al., 2023; Together.ai, 2023), 128K (OpenAI, 2023; Fu et al., 2025), or even millions of tokens (AI, 2024; Cloud, 2024; Liu et al., 2024b).
+Considering the extrapolation problem (Press et al., 2022; Liu et al., 2024a), which refers to the decline in LLM performance when input sequences exceed the training length, it is necessary to conduct long context training (Jacobs et al., 2023; Chen et al., 2024; Fang and Zhao, 2024) or fine-tuning (Peng et al., 2024; Ding et al., 2024) to facilitate long sequence inference.
 Beyond natural language processing, increasing the context length is also essential across diverse domains, including video processing (Zheng et al., 2024), protein properties prediction (Chandra et al., 2023), weather forecasting (Nguyen et al., 2023), and health care (Li et al., 2022).
-However, training LLMs with long sequence lengths poses a significant challenge for GPU memory.
-During training, a large amount of activations (i.e., the intermediate results computed in the forward pass) must be stored for gradient computation during the backward pass, resulting in substantial memory consumption.
+Maximizing system performance with limited memory is a common and significant challenge in the data management community.
+Within this context, training LLMs with long sequence lengths poses difficulties due to restricted GPU memory.
+During training, a large amount of activations [^1] In neural network training, the outputs of operators are referred to as activations.
+Some of these, termed “skeletal activations” in this work, must be stored for gradient computation during the backward pass, while others are termed “transient activations”.
+More details are provided in Section 3.
+must be stored for gradient computation during the backward pass, resulting in substantial memory consumption.
 Typically, it is well known that the self-attention module in the transformer architecture has a quadratic computation and memory complexity w.r.t.
 the sequence length.
-FlashAttention (Dao et al., 2022; Dao, 2023), now a standard technique for attention computation in LLM training, accelerates computation and shrinks the memory complexity to be linear w.r.t.
+FlashAttention (Dao et al., 2022; Dao, 2024), now a standard technique for attention computation in LLM training, accelerates computation and shrinks the memory complexity to be linear w.r.t.
 the sequence length by scheduling memory I/O and recomputing necessary components during the backward pass.
-Except for attention, the remaining activation memory also scales proportionally with the sequence length, which can become quite large in long context scenarios.
+Except for attention, the remaining activation memory also scales linearly with the sequence length, which can become quite large in long context scenarios.
 For instance, training a GPT model with 7B parameters on a sequence length of 1 million can lead to an activation memory of 4096GB, far exceeding the memory capacity of commonly used accelerators (e.g.
 80GB for an NVIDIA H100/A100 GPU).
-Moreover, the memory fragmentation issue makes this situation even worse.
+Moreover, dynamic memory allocators inherently face the issue of memory fragmentation due to frequent allocation and deallocation, which complicates efficient data management.
 Besides storing the skeletal activations for the backward pass, there are also tremendous transient activations that are temporarily generated during computation (we will formally categorize the two kinds of activations in Section 3).
@@
 This issue becomes more severe in long context training, considering the fact that the (de)allocated memory blocks are significantly larger than those in normal tasks.
-Memory fragmentation not only leads to Out-Of-Memory (OOM) error but also significantly hinders training efficiency because of the frequent triggering of the PyTorch memory reorganization process, which involves calls to expensive “cudaFree” and “cudaMalloc” to release cached blocks and reclaim GPU memory.
-Figure 1(a) illustrates an example of GPU memory fragmentation.
+Memory fragmentation not only leads to Out-Of-Memory (OOM) error but also significantly hinders training efficiency because of the frequent triggering of the PyTorch memory reorganization process, which involves calls to “cudaFree” and “cudaMalloc” to release cached blocks and reclaim GPU memory.
+Figure 1 illustrates an example of GPU memory fragmentation.
 At the peaks of the curves, there is more than 4GB memory reserved but not allocated.
 However, when the training task tries to allocates 4GB memory, the allocator fails to find a continuous memory space to fulfill the allocation request.
-Consequently, it necessitates to invoke a series of “cudaFree” and “cudaMalloc” to reorganize memory, which blocks GPU computation.
-(a) GPU memory consumption.
-(b) One-layer forward time consumption.
+Consequently, it necessitates invoking a series of “cudaFree” and “cudaMalloc” to reorganize memory, which blocks GPU computation.
 Figure 1.
@@
 There are several key observations that inspire our design.
-Observation 1: Opportunity for activation swapping.
-In the field of deep learning training, to reduce the peak memory consumption caused by skeletal activations, activation recomputation  (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) and swapping (Rhu et al., 2016; NVIDIA, 2024b) are two well-known memory reduction techniques that trade time to save memory.
-[^1] Parallelism techniques like sequence parallelism (Korthikanti et al., 2022; Jacobs et al., 2023) and context parallelism (Liu et al., 2023b; Li et al., 2023; NVIDIA, 2024a) are also compelling approaches to reduce memory at the price of extra communication overhead.
+In the data management domain, when high-bandwidth memory is limited, rematerialization is a typical technique to free up memory by releasing data that is not immediately needed and reconstructing data structures on demand.
+For instance, Spark (Zaharia et al., 2016) supports rematerializing Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) through recomputation or swapping [^2] Swapping refers to moving data between different levels of memory hierarchy.
+More details are provided in Section $2.$ between CPU memory and disk storage when RAM is insufficient.
+Faiss (Johnson et al., 2021), a widely used vector database, utilizes a CPU-GPU memory hierarchy to accommodate the high memory requirement.
+Additionally, the dynamic view materialization solution (Phan and Li, 2008) employs an LRU cache to manage materialized views, and rematerializes cache-missed views upon needed.
+These methods leverage recomputation and swapping strategies to efficiently manage data structures within the memory hierarchy.
+Similarly, in deep learning training, to reduce the peak memory consumption caused by skeletal activations, activation recomputation  (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) and swapping (Rhu et al., 2016; NVIDIA, 2024b) are also widely adopted.
+[^3] Parallelism techniques like sequence parallelism (Korthikanti et al., 2022; Jacobs et al., 2023) and context parallelism (Liu et al., 2023; Li et al., 2023a; NVIDIA, 2024a) are also compelling approaches to reduce memory at the price of extra communication overhead.
 Our work is compatible with these parallelism techniques.
 Typically, both of them reduce memory consumption at the price of extra time cost.
-For one thing, the activation recomputation technique discards the skeletal activations in the forward pass and later recomputes them in the backward pass, leading to extra computation cost.
-For another, the swapping technique offloads the activations to CPU memory in the forward pass to relieve the GPU memory pressure, and later fetches them back to GPU memory in the backward pass, incurring the overhead of data transmission between CPU and GPU memory.
-Contemporary, mainstream LLM training frameworks such as Megatron-LM and DeepSpeed prefer activation recomputation to swapping, which is due to the fact that the GPU computing ability has a far more rapid growth than the connectivity between CPU and GPU memory in the past few years (see Section 2.2 for details).
+The activation recomputation technique discards some activations in the forward pass and later recomputes them in the backward pass, leading to extra computation cost.
+The swapping technique offloads the activations to CPU memory in the forward pass to relieve the GPU memory pressure, and later fetches them back to GPU memory in the backward pass, incurring the overhead of data transmission between CPU and GPU memory.
+Observation 1: Opportunity for activation swapping.
+Contemporary mainstream LLM training frameworks such as Megatron-LM and DeepSpeed prefer activation recomputation to swapping, which is due to the fact that the GPU computing ability has a far more rapid growth than the connectivity between CPU and GPU memory in the past few years (see Section 2.2 for details).
 However, we find that the situation is a bit different in long context training of LLMs.
@@
 As the sequence length increases, there is greater potential for overlapping computation and communication, given that their time requirements scale quadratically and linearly with the sequence length, respectively.
-As shown in Figure 1(b), eventually, after reaching a specific sequence length (192K in this case), the transmission of activations can be fully overlapped with GPU computation.
-However, in practice, there is limited chance to completely swap all activations.
+As shown in Figure 1, eventually, after reaching a specific sequence length (192K in this case), the transmission of activations can be fully overlapped with GPU computation.
+Figure 2.
+An overview of Memo.
+We devise a fine-grained recomputation and swapping mechanism to manage the skeletal activations for backward propagation, and leverage a bi-level memory planning method to reuse the memory space of transient activations across the transformer layers.
+Memo Overview
+However, in practice, there is limited chance to swap all activations.
 On the one hand, extremely long training data is rare, and most of the time we need to train on data that doesn’t fully overlap the activation transmission and the computation.
@@
 We consider both tensor-level and token-level activation management.
-For each layer, following previous works (NVIDIA, 2024b; Li et al., 2023), we consistently offload two kinds of activation tensors, the input of each transformer layer and the output of FlashAttention, to CPU memory.
+For each layer, following previous works (NVIDIA, 2024b; Li et al., 2023a), we consistently offload two activation tensors, the input of each transformer layer and the output of FlashAttention, to CPU memory.
 For other activation tensors, we only offload a fraction (denoted as $\alpha$) of tokens, and recompute the rest part during the backward pass.
-We model the time cost of activation recomputation and transmission and determine the fraction $\alpha$ through a well-formulated linear programmin problem, which aims to maximize offloading activations without impeding GPU computation or causing CPU OOM issues.
+We model the time cost of activation recomputation and transmission and determine the fraction $\alpha$ through a well-formulated linear programming problem, which aims to maximize offloading activations without impeding GPU computation or causing CPU OOM issues.
 During the backward pass, prefetching activations can also overlap with GPU computation, because the backward computation is typically twice as much as the forward computation.
 With both tensor-level and token-level activation management, we make full use of the idle bandwidth and minimize the recomputation overhead to improve the overall efficiency.
-Figure 2.
-An overview of Memo.
-We devise a fine-grained recomputation and swapping mechanism to manage the skeletal activations for backward propagation, and leverage a bi-level memory planning to reuse the memory space of transient activations across the transformer layers.
-Memo Overview
+Regarding memory fragmentation, research has utilized the characteristics of targeted workloads to analyze and resolve the issue, such as experimentally analyzing the impact of memory allocation on high-performance query engines (Durner et al., 2019) and addressing persistent memory fragmentation with efficient defragmentation algorithms (Oukid et al., 2017).
+In the same vein, for long-sequence LLM training, we aim to leverage the specific characteristics of LLM training to address memory fragmentation.
 Observation 2: Deterministic memory (de)allocation pattern across iterations and layers.
-The memory fragmentation mainly comes from frequent and irregular memory (de)allocation requests.
+In long sequence LLM training, the memory fragmentation mainly comes from frequent and irregular memory (de)allocation requests.
 However, we observe that, typical LLM training adheres to a deterministic computation process across iterations and layers.
@@
 Before training, we profile the memory (de)allocation requests of one training iteration, then use MIP to solve an optimized memory plan for a single transformer layer.
-Since each layer’s memory request are identical, the entire memory block for one layer can be directly reused for the subsequent identical layer.
+Since the memory requests of the transformer layers are identical, the entire memory block for one layer can be directly reused for the subsequent identical layer.
 Considering each transformer layer’s memory block as a single memory allocation request, we further solve another MIP problem that plans memory allocation for the entire LLM training, including the initial embedding layer, all transformer layers, and the final classifier layer.
@@
 Putting them together, in response to the activation memory challenge in long context training, we propose Memo, an LLM training framework with fine-grained tensor memory management.
-We consider the challenge as an activation data management problem.
+We consider the challenge as an activation data management problem, and draw inspiration from the data rematerialization and memory defragmentation techniques to address the challenge.
+Figure 2 presents an overview of Memo.
 To make full use of the idle CPU-GPU bandwidth during training with different sequence lengths, we introduce a token-wise fine-grained activation recomputation and swapping strategy.
@@
   • We introduce a fine-grained activation recomputation and swapping mechanism to fully utilize the idle CPU-GPU communication bandwidth during time-consuming GPU computation.
-  • We employ a bi-level hierarchical MIP technique to solve the memory planning problem and significantly mitigate memory fragmentation.
-  • We evaluate Memo through extensive experiments, and demonstrate an average of $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
+  • We employ a bi-level MIP technique to solve the memory planning problem and significantly mitigate memory fragmentation.
+  • We evaluate Memo through extensive experiments, and demonstrate an average of $1.97\times$ and $1.80\times$ improvement in terms of MFU [^4] MFU (Model FLOPs Utilization) is a widely used efficiency metric that evaluates how well the accelerators are utilized in model training.
+It is calculated as the ratio of the observed throughput to the theoretical throughput which assumes the hardware operates at peak FLoating-point Operations Per Second (FLOPS) (Pope et al., 2023) .
+More details are provided in Section 5.1.
+compared to Megatron-LM and DeepSpeed, respectively.
 Additionally, Memo is the first framework that enables the efficient training of 7B LLM with 1 million context length on only 8 A800 GPUs.
@@
 Commonly used notations are listed in Table 1.
-Figure 3. The architecture of a typical LLM.
-gpt_arch
+(a) The architecture of a typical LLM.
+(b) An example memory request sequence.
+Figure 3.
+(a): The architecture of a typical LLM.
+(b): An example memory request sequence of a transformer layer’s forward and backward pass.
+Tensors 15 and 16 are skeletal tensors, while the others are transient tensors.
 2.1. Large Language Models
 2.1.1. Architecture
-As shown in Figure 3, the architecture of a typical LLM comprises an input embedding layer, multiple decoder-only transformer layers, and a final classifier layer.
+As shown in Figure 3(a), the architecture of an LLM comprises an input embedding layer, multiple transformer layers, and a final classifier layer.
 The embedding layer converts input tokens into continuous representations.
@@
 2.1.3. The Challenge of Huge Memory Requirement in Long Context Training
-In the original form of self-attention, which is the most critical module in LLMs, input tokens are initially projected into queries, keys, and values.
-The queries and keys are multiplied and then softmaxed to create an $s\times s$ matrix that represents attention weights, which are subsequently used to compute a weighted sum of the values.
-Thus, it exhibits a memory complexity of $O(s^{2})$.
-As the sequence length increases, storing the entire $s\times s$ tensors for the backward pass becomes infeasible.
-FlashAttention (Dao et al., 2022; Dao, 2023) has emerged as a solution for efficient attention computation in long context training.
-It minimizes memory I/O and carries out the attention computation in a streaming fashion in both the forward and backward pass, and thereby successfully get rid of the huge $s\times s$ matrix.
-With such designs, FlashAttention accelerates the overall computation speed (while maintaining a time complexity of $O(s^{2})$), and eliminates the $O(s^{2})$ memory requirement.
-Currently, FlashAttention has become the de-facto strategy for self-attention computation.
-Thus, we assume FlashAttention is employed throughout this work.
+Self-attention is the most critical module in LLMs.
+It facilitates information interaction between tokens: the input tensor is first projected into query ($Q$), key ($K$), and value ($V$), each with the shape ($b,s,h$); then the output is given by $O=\text{softmax}({QK^{T}}/{\sqrt{d}})\cdot V$, which incurs $O(s^{2})$ time and space complexity due to the ($b,s,s$) matrices.
+FlashAttention (Dao et al., 2022; Dao, 2024), the de-facto attention implementation in nowadays LLM computation, processes the computation in tiles, discards intermediate results, and maintains compact states to generate the final output $O$.
+This method avoids storing the $O(s^{2})$ matrices.
+During the backward pass, FlashAttention re-computes these intermediate results in a tiled manner for gradient calculation.
+Thanks to this design, FlashAttention significantly reduces memory requirements to just $O(s)$ complexity.
+Although several alternative approaches like sparsification (Miao et al., 2024b) (e.g., BigBird (Zaheer et al., 2020), Longformer (Beltagy et al., 2020)), kernelization (e.g., LinearAttn (Katharopoulos et al., 2020), CosFormer (Qin et al., 2022)), and low-rank approximation (e.g., Linformer (Wang et al., 2020), Nyströmformer (Xiong et al., 2021)) also aim to reduce the quadratic memory demands of self-attention, these methods modify the attention mechanism and can potentially compromise accuracy.
+In this paper, we adopt FlashAttention as the default method, considering its prevalent use in practice.
 Although FlashAttention has reduced the memory complexity of LLM training from $O(s^{2})$ to $O(s)$, the linearly scaling activation memory remains the primary challenge in long context training.
@@
 It is worth noting that although our work primarily concentrates on the memory reduction techniques, the proposed Memo framework is compatible with a wide range of parallelism strategies.
-2.2. Memory Reduction Techniques
-Limited GPU memory has become the bottleneck of long context LLM training.
-To alleviate the memory pressure, there are two notable memory reduction techniques, namely activation recomputation and swapping.
+$2.$ Memory Reduction Techniques
+As mentioned in Section 1, when existing query processing engines face the challenge of limited high-bandwidth memory, a common strategy to mitigate memory pressure involves discarding certain data structures and rematerializing them as needed (Zaharia et al., 2016; Johnson et al., 2021; Phan and Li, 2008).
+There are two prevalent techniques: (1) recomputing the results and (2) swapping data to lower-tier memory and retrieving it when necessary.
+Both methods are also widely-used in neural network training to rematerialize activation tensors.
 Activation recomputation (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) (a.k.a.
-activation checkpointing) selectively stores only the inputs of certain layers rather than storing all intermediate activations.
+activation checkpointing) selectively stores the inputs of certain layers rather than all intermediate activations.
 During the backward pass, the required activations are recomputed on-the-fly.
@@
 Swapping (Rhu et al., 2016; NVIDIA, 2024b; Ren et al., 2021), also known as CPU offloading, aims to relieve the GPU memory pressure by offloading GPU tensors to CPU memory, and fetch them back to GPU when needed.
-Although swapping does not consumes GPU computation units, it would still slow down the training if the data transmission cannot be overlapped by GPU computation.
-In general, both the two memory reduction techniques release the memory of activations in the forward pass, but need to rematerialize them in the backward pass, at the price of extra computation or data transmission overhead, respectively.
+Through careful scheduling, the data transmission overhead can be overlapped with GPU computation, a technique also popular in GPU databases (Johnson et al., 2021; Rui et al., 2020; Wang et al., 2014).
+However, if data transmission is too time-consuming to overlap, swapping can significantly slow down training.
+In general, both memory reduction techniques release the memory of activations in the forward pass, but need to rematerialize them in the backward pass, at the price of extra computation or data transmission overhead, respectively.
 […]
 2.3. Distributed Parallelism Strategies
-Distributed training is essential for efficiently training LLMs, especially in the senario of long context training.
+Distributed training is essential for efficiently training LLMs, especially in scenarios of long context training.
 To facilitate the training of large-scale data and model, several distributed parallelism strategies have been proposed.
@@
 Zero-Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2019) is a series of variants built upon DP, aiming to alleviate memory pressure.
-In naive DP, model parameters, gradients and optimizer states are replicated among all devices.
+Naive DP replicates model parameters, gradients and optimizer states among all devices.
 ZeRO is designed in three stages to reduce these memory requirements respectively.
@@
 Next, ZeRO-2 extends ZeRO-1 by also partitioning gradients, further reducing memory footprint.
-Finally, ZeRO-3, based on ZeRO-2, partitions model parameters among DP workers, further mitigating memory pressure but introducing additional communication to gather parameters during forward and backward pass.
+Finally, ZeRO-3, based on ZeRO-2, partitions model parameters among DP workers, further mitigating memory pressure but introducing additional communication to gather parameters during training.
 Tensor Parallelism (TP) (Shoeybi et al., 2019) partitions the self-attention and feed-forward modules of transformer layers across multiple devices along either the column or row dimension.
@@
 Therefore, TP is usually applied within a computing node, where intra-node GPUs are connected via high-bandwidth NVLink.
-Pipeline Parallelism (PP) (Huang et al., 2019; Narayanan et al., 2021) is also proposed to address the problem that LLMs cannot be fit onto a single device.
-Different from TP, PP partitions different model layers into several stages, then distributes the stages to different devices.
+Pipeline Parallelism ($P$) (Huang et al., 2019; Narayanan et al., 2021; Guan et al., 2024) is also proposed to address the problem that LLMs cannot be fit into a single device.
+Different from TP, $P$ partitions model layers into several stages, then distributes the stages to different devices.
 The input data is processed through these stages in a pipeline fashion.
-Given the peer-to-peer communication style, the PP stages are often distributed across nodes.
-However, PP introduces a phenomenon known as “bubble”, which corresponds to GPU idle time.
+Given the peer-to-peer communication style, the $P$ stages are often distributed across nodes.
+However, $P$ introduces a phenomenon known as “bubble”, which corresponds to GPU idle time.
 The issue becomes more severe when the number of micro-batches is small.
-To facilitate efficient long context training, several novel parallelism strategies have been recently proposed.
+To facilitate efficient long context training, several novel parallelism strategies have been proposed recently.
 Sequence Parallelism (SP) (Korthikanti et al., 2022) is built upon TP to further reduce activation memory overhead.
 It splits the sequence dimension in the part of the model that does not apply TP.
-The original AllReduce communication now transitions to AllGather and ReduceScatter communication.
+The original AllReduce communication now transitions to AllGather and ReduceScatter.
 DeepSpeed-Ulysses (Jacobs et al., 2023), built upon ZeRO, is another form of sequence parallelism.
@@
 To further relieve the memory pressure, DeepSpeed-Ulysses leverages ZeRO to distribute model parameters.
-Context Parallelism (CP) (Liu et al., 2023b; Li et al., 2023; NVIDIA, 2024a) proposes sharding the query, key, value matrices within the self-attention module along the sequence dimension across different devices.
-During attention computation, necessary communications are involved to ensure consistent results.
-The communications can be overlapped with computation by careful scheduling.
+Context Parallelism (CP) (Liu et al., 2023; Li et al., 2023a; NVIDIA, 2024a) shards the query, key, and value matrices within the attention module along the sequence dimension across different devices.
+During attention computation, necessary communications are involved to ensure consistent results, which can be overlapped with computation by careful scheduling.
 In practice, these parallelism strategies and memory reduction techniques can be integrated and employed simultaneously to facilitate efficient training of LLMs.
-Figure 4. An example memory request sequence of a transformer layer’s forward and backward propagation. Tensor 15 and 16 are skeletal tensors, while the others are transient tensors.
-An example memory sequence
 3. Anatomy and System Desiderata
-Given the fact that the memory consumption of activations scales proportionally w.r.t.
-the sequence length in LLM training, we first provide an in-depth anatomy of the key characteristics of different activations in this section.
-Based on this analysis, we present the design desiderata that motivate the development of Memo.
+Managing fragmented massive data storage is a critical data management issue, particularly when workloads are constrained by limited high-bandwidth memory.
+Long-context LLM training demonstrates such a challenge, where huge fragmented activation memory significantly impedes efficient training within constrained GPU memory resources.
+In this section, we first provide an in-depth anatomy of the key characteristics of activation data storage in long-context LLM training.
+Based on this analysis, we present the design desiderata that motivates the development of Memo.
 3.1. Categorization of Activation Tensors
-To be specific, according to their life cycles, we categorize activations generated during the forward propagation into two classes, which are the skeletal activations and the transient activations, where the former is necessary for the backward propagation while the latter is not.
-For illustration, in Figure 4, tensors 13, 14, 17, 18, and 19 are produced during the forward propagation of a transformer layer, and are discarded before the completion of this layer’s forward pass.
-Similarly, tensors 20, 21, 22, 23, and 24 are generated during the backward propagation of this layer, and are discarded after corresponding computation.
-We term them “transient tensors” because they are both created and discarded within a single layer’s forward or backward pass.
+In long-context LLM training, the primary memory consumption originates from activation tensors, which are the outputs of computing operators in LLMs.
+According to their life cycles, we can categorize activations generated during the forward propagation into two classes, which are the skeletal activations and the transient activations, where the former is necessary for the backward propagation while the latter is not.
+For illustration, in Figure 3(b), tensors 13, 14, 17, 18, and 19 are produced during the forward pass of a transformer layer, and are discarded before the completion of this layer’s forward pass.
+Similarly, tensors 20, 21, 22, 23, and 24 are generated during the backward pass of this layer, and are discarded after corresponding computation.
+We term them “transient tensors” because they are created and discarded within a single layer’s forward or backward pass.
 Transient tensors usually serve as temporary results.
@@
 Forward and backward propagation with rounding buffers for token-wise recomputation/swapping.
-During forward propagation, the darker part in the rounding buffers is offloaded to CPU, while the lighter part is discarded; during backward propagation, the darker part in the rounding buffers if prefetched from CPU, while the lighter part it recomputed.
+During forward propagation, the darker part in the rounding buffers is offloaded to CPU, while the lighter part is discarded; during backward propagation, the darker part in the rounding buffers is prefetched from CPU, while the lighter part is recomputed.
 Rounding buffers with token-wise recomputation
@@
 To exemplify, when training the GPT-7B model ($h=4096$, 32 layers) with a sequence length ($s$) of 1 million, if we store the skeletal activations in half-precision floating numbers, it would take 4096 GB for only one sequence ($b=1$), exceeding the memory capacity of even 50 A100/H100 GPUs.
-The most important characteristic of skeletal activations is that they are needed by backward computation, so they must reside in GPU memory at least before the backward propagation of the corresponding transformer layer begins.
-However, there is no doubt that maintaining all skeletal activations for backward propagation is infeasible.
-To this end, memory-saving techniques like activation recomputation and swapping become necessary for long context training.
-These techniques first release the skeletal activations of a transformer layer in the forward propagation, and later rematerialize them before the corresponding backward propagation.
-In essence, both the two techniques trade time for memory — the activation recomputation technique incurs extra computation overhead while the swapping technique necessitates transmitting the activations from CPU memory to GPU memory.
-For extremely long context lengths, using either technique would take significant time to rematerialize the skeletal activations, causing performance degradation.
+An important characteristic of skeletal activations is that they are needed by backward computation, so they must reside in GPU at least before the backward propagation of the corresponding transformer layer begins.
+However, maintaining all skeletal activations for backward propagation is infeasible.
+To this end, memory-saving techniques like recomputation and swapping, which are also prevalent in traditional data management problems (Zaharia et al., 2016; Johnson et al., 2021; Phan and Li, 2008), become necessary for long context training.
+In LLM training, these techniques first release the skeletal activations of a transformer layer in the forward propagation, and later rematerialize them before the corresponding backward propagation.
+Unfortunately, naïvely applying activation recomputation or swapping is insufficient to tackle the challenge of managing large-scale skeletal activations.
+Both techniques trade time for memory — the activation recomputation technique incurs extra computation overhead while the swapping technique necessitates transmitting the activations from CPU memory to GPU memory.
+Using recomputation alone incurs significant additional computation overhead, while employing swapping alone can lead to CPU OOM error (when the sequence is too long) or block GPU computation (when the swapping time cannot be fully overlapped).
 As a result, we desiderate a meticulous orchestration of the two memory-saving techniques to manage the skeletal activations, so that we can minimize the extra overhead while accommodating the huge memory requirement in long context training of LLMs.
-To achieve this goal, we develop a token-wise activation recomputation and swapping mechanism, which will be demonstrated in Section 4.1.
-3.3. Analysis of Transient Activations
+To achieve this, we develop a token-wise activation recomputation and swapping mechanism, which will be demonstrated in Section 4.1.
+$3.$ Analysis of Transient Activations
 Transient activations are intermediate results generated and discarded during the forward (or backward) pass of a transformer layer.
@@
 Specifically, we observe that the number of transient activations can exceed 5 times that of skeletal activations.
-Without careful management, the frequent allocation and deallocation can lead to memory fragmentation and degrades system performance.
-Fortunately, the most important characteristic of transient activations is that they are identical cross transformer layers, which provides us with the opportunity to manage and reuse their memory regions to minimize fragmentation.
+Without careful management, the frequent allocation and deallocation can lead to memory fragmentation, which degrades system performance and poses a major concern for managing massive data storage.
+There have been studies in the data management community focusing on memory defragmentation (Durner et al., 2019; Oukid et al., 2017), leveraging the characteristics of target workloads to devise appropriate and innovative methods.
+Following the methodology, we analyze the LLM training process and attempt to defragment the tensor memory.
+During training, memory requests are identical across both transformer layers and training iterations, providing an opportunity to manage and reuse these memory regions effectively to minimize fragmentation.
 In particular, the memory addresses of a single transformer layer’s transient activation tensors can be reused by all other transformer layer’s corresponding transient activation tensors.
@@
 sequence length.
-And we propose token-wise activation recomputation and swapping, along with a bi-level memory planning to address the issue, which targets skeletal activation tensors and transient activation tensors, respectively.
+We propose token-wise activation recomputation and swapping, along with a bi-level memory planning to address the issue, which targets skeletal activations and transient activations, respectively.
 The overview of Memo is depicted in Figure 2.
@@
 In practice, as sequence length grows, the size of skeletal activations increases linearly, which can easily exceed the capacity of GPU memory.
-As introduced in Section 2.2, currently the most widely-used technique to tackle this issue is activation recomputation, which stores only the input of each transformer layer, and discard the rest skeletal activation tensors of this layer.
+As introduced in Section 2.2, currently the most widely-used technique to tackle this issue is activation recomputation, which stores only the input of each transformer layer, and discards the rest skeletal activation tensors of this layer.
 Prior to backward propagation of each layer, an additional forward pass of the layer is conducted to reconstruct all skeletal tensors so that the backward computation can be carried out.
@@
 The prefetching of transformer layer $i$’s skeletal activations happens simultaneously with the backward propagation of transformer layer $(i+1)$.
-When the sequence length is sufficiently long, with careful computation-transmission overlapping and synchronization, CPU swapping can substitute activation recomputation without incurring additional computation overhead.
+When the sequence length is sufficiently long, with careful computation-transmission overlapping and synchronization, CPU swapping can substitute activation recomputation without incurring additional overhead.
 However, there are two constraints that prevent us from offloading all skeletal activations to CPU memory.
@@
 This discrepancy forces the computation of transformer layer $(i+2)$ to be delayed until the offloading of rounding buffer to CPU memory is completed, thereby blocks the normal GPU computation workflow.
-For instance, as illustrated in Figure 1(b), when training a 7B GPT model on 8 GPUs with a TP size of 8, ideal overlap between the computation of a transformer layer and the offloading of its skeletal activations occurs only for sequence lengths exceeding 192K.
-In practice, the sequence lengths of most training datasets for LLMs are moderate and may be not sufficient to ensure an ideal overlap between computation and transmission.
-  • In theory, a longer sequence length provides more opportunities for overlapping CPU offloading with GPU computation.
-However, in practice, the CPU memory in GPU servers is often limited.
+For instance, as illustrated in Figure 1, when training a 7B GPT model on 8 GPUs with a TP degree of 8, ideal overlap between the computation of a transformer layer and the offloading of its skeletal activations occurs only for sequence lengths exceeding 192K.
+In practice, the sequence lengths of most training datasets are moderate and may be not sufficient to ensure an ideal overlap between computation and transmission.
+  • In theory, a longer sequence length provides more opportunities for overlapping swapping with GPU computation.
+However, in practice, the CPU memory is often limited.
 For a typical GPU server which has several terabytes CPU memory (e.g.
-2TB in our environment), this is insufficient to store all skeletal activations when the sequence length is excessively long or the number of transformer layers is too large.
+2TB in our environment), it is insufficient to store all skeletal activations when the sequence length is excessively long or the number of transformer layers is too large.
 For instance, when training the 7B model on a server equipped with 8 GPUs using a sequence length of 1 million, the skeletal activations amount to a total size of 4096GB, which is double the capacity of CPU memory.
-Figure 7. Forward time of FlashAttention and other parts of a transformer layer when training a 7B GPT on 8 GPUs with a TP size of 8.
+Figure 6. Forward time of FlashAttention and other parts of a transformer layer when training a 7B GPT on 8 GPUs with a TP degree of 8.
 Forward time of FlashAttention and other parts of a transformer layer
 Therefore, instead of simply offloading all skeletal activations to CPU memory, we employ selective activation swapping to ensure perfect overlap of computation and transmission for short sequences as well as to avoid depleting CPU memory for extremely long context lengths.
-Memo manages to determine the selection of swapping at both the tensor and token granularities, as depicted in Figure 6.
+Memo determines the selection of swapping at both the tensor and token granularities, as depicted in Figure 5.
 At the tensor granularity, we consider the benefits of leveraging the swapping technique rather than the recomputation technique of different modules.
@@
 Besides, since LLMs have a layered structure, in order to reconstruct the “input_norm”, “q”, “k”, “v” tensors, we also store the input of each transformer layer to CPU, following common recomputation strategy (Chen et al., 2016).
-At the token granularity, we develop the token-wise activation recomputation and swapping technique to reduce the memory consumption of all skeletal activation tensors other than the output of FlashAttention and the input of each transformer layer.
-To be specific, as shown in Figure 6, for each of these skeletal activation tensors, we only offload a fraction (denoted as $\alpha$) to CPU, while the remaining part is discarded, ensuring perfect overlapping and to avoid CPU OOM error.
-Before the backward propagation, the discarded part is rematerialized via recomputation while the offloaded part is prefetched.
+At the token granularity, we develop the token-wise activation recomputation and swapping technique to reduce the memory consumption of all skeletal activation tensors other than the output of FlashAttention and the input of each layer.
+To be specific, as shown in Figure 5, for each of these skeletal activation tensors, we offload a fraction (denoted as $\alpha$) to CPU, while the remaining part is discarded, ensuring perfect overlapping and to avoid CPU OOM error.
+Before the backward pass, the discarded part is rematerialized via recomputation while the offloaded part is prefetched.
 To determine the fraction $\alpha$, we solve the following problem:
-(1) $\max\quad$ $\alpha,$
-(2) s.t. $(S_{input}+S_{attn}+\alpha\cdot S_{others})/B\leq T_{layer},$
-(3) $(n-2)(S_{input}+S_{attn}+\alpha\cdot S_{others})\leq M_{CPU}.$
-where $S_{input}$, $S_{attn}$, and $S_{other}$ stand for the size of input tensor, the size of FlashAttention output tensor, the total size of other skeletal activation tensors, respectively, $B$ is the PCIe bandwidth between GPU and CPU, $T_{layer}$ is the forward time of a single transformer layer, $n$ is the total number of transformer layers, and $M_{CPU}$ stands for the capacity of CPU memory.
+$\max\quad$ $\alpha,$
+s.t. $(S_{input}+S_{attn}+\alpha\cdot S_{others})/B\leq T_{layer},$
+$(n-2)(S_{input}+S_{attn}+\alpha\cdot S_{others})\leq M_{CPU}.$
+where $S_{input}$, $S_{attn}$, and $S_{other}$ stand for the size of input tensor, the size of FlashAttention output tensor, the total size of other skeletal activation tensors, respectively; $B$ is the PCIe bandwidth between GPU and CPU, $T_{layer}$ is the forward time of a single transformer layer, $n$ is the total number of transformer layers, and $M_{CPU}$ stands for the capacity of CPU memory.
 It is worth noting that, the last two transformer layers can initiate the backward pass immediately after the forward pass, obviating the need for swapping.
 These variables can be easily obtained through profiling before training, so we can determine an appropriate $\alpha$ without much effort.
-As a special case, when the determined fraction is 0, skeletal buffers except layer input and attention output can share one GPU buffer, as they are fully recomputed and do not need to be offloaded.
-Thus there is no need to use two rounding buffers to avoid data corruption.
 4.2. Bi-level Memory Planning
@@
 A concise overview of this formulation is shown as follows.
-The offline DSA problem assumes a predefined sequence of memory allocations and deallocations, and aims to determine the address of each allocated memory block and at the same time minimizing the peak memory usage.
+The offline DSA problem handles a sequence of memory allocations and deallocations, and aims to determine the address of each allocated memory block and at the same time minimizing the peak memory usage.
 Parameters of offline DSA problem includes:
@@
 Following this formulation, the solution for each tensor’s address is optimal.
-However, modern LLM training involves thousands of allocation and deallocation requests, which makes this MIP problem computationally intractable.
+However, modern LLM training involves thousands of allocation and deallocation requests within a single training iteration, which makes this NP-hard MIP problem computationally intractable.
+Consequently, it’s infeasible to solve this MIP problem in one pass, given the prohibitively high time cost.
 Fortunately, all transformer layers have identical structures and memory request sequences, which presents repetitive substructures within the MIP problem.
-By leveraging this inherent repetitiveness, we can devise a bi-level hierarchical MIP optimization algorithm, which is both computationally feasible and effective.
+By leveraging this inherent repetitiveness, we can instead devise a bi-level hierarchical MIP optimization algorithm, which is both computationally feasible and effective.
 Figure 8. Memory request sequence during training.
@@
 As shown in the bottom of Figure 7, we first solve the offline DSA sub-problem for just one transformer layer’s forward (backward) pass, which is called the first-level MIP.
-This offline DSA problem can be simply solved by any MIP solver.
+The scale of the first-level MIP is small enough to tackle.
+This offline DSA problem can be simply solved by any MIP solver (e.g.
+Gurobi (Gurobi Optimization, LLC, 2024)).
 After this step, the peak memory needed for the forward (backward) propagation of a single transformer layer, as well as the address of each transient tensor within a transformer layer is determined.
 After solving the sub-problem for one transformer layer, all other transformer layers can reuse the same memory address for (de)allocation.
-Subsequently, we can replace the original fine-grained memory request sequence of a transformer layer forward (backward) propagation with a “pseudo” large memory request pair, as shown in Figure 8.
+Besides identical transformer layers, LLMs also have other layers that process input tokens and classify output tokens.
+Therefore, after solving the first MIP for one-layer DSA problem, we conduct the second MIP for the whole LLM training to generate the peak memory requirement and addresses of all transient activation tensors.
+To simplify the optimization process, we can replace the original fine-grained memory request sequence of each transformer layer’s forward (backward) propagation with a “pseudo” large memory request pair, as shown in Figure 7.
+The size of this “pseudo” memory block corresponds to the memory usage of each transformer layer as determined by the first-level MIP.
 After the substitution, this reformulated memory request sequence also satisfies the formulation of an offline DSA problem, with a size small enough to be efficiently solved.
 We then leverage the MIP solver again to solve this second-level MIP problem.
-After this step, the addresses of all activation tensors, as well as the peak memory needed for all transient activation tensors can be determined.
+After this step, the addresses of all activation tensors, and the peak memory needed for all transient activation tensors can be determined.
 4.3. System Implementation
@@
 Figure 9 illustrates the overall architecture of Memo.
-First, the job profiler takes in the model configuration and user-defined settings, then executes a training iteration to profile the memory requests directed to the PyTorch CUDA allocator during the training phase.
+First, the job profiler takes in the model configuration, then executes a training iteration to profile the memory requests directed to the PyTorch CUDA allocator during the training phase.
 The job profiler also determines offloading fraction $\alpha$ by solving optimization problem in Section 4.1.
 These memory requests comprise a sequence of allocation and deallocation instructions.
-Afterwards, the memory planner receives the memory requests, executes the bi-level MIP optimization algorithm and, generates a memory plan, which constitutes the addresses of all activation tensors during one training iteration.
+Afterwards, the memory planner receives the memory requests, executes the bi-level MIP optimization algorithm and, generates a memory plan, which constitutes the addresses of all transient activation tensors during one training iteration.
 Finally, the runtime executor reads the memory plan and conducts the training process.
@@
 However, naively recording all memory requests may lead to OOM error.
-For example, profiling a GPT-7B model with a sequence length of 512K on 8 GPUs can result in OOM error since the proposed techniques are not employed in profiling.
-Fortunately, as the transformer layers have identical memory footprint, we leverage this property by only profiling a single transformer layer’s memory footprint and then applying it to all transformer layers.
+For example, directly profiling a GPT-7B model with a sequence length of 512K on 8 GPUs can result in OOM error.
+Fortunately, all transformer layers have identical memory footprint.
+We leverage this property by only profiling one transformer layer’s memory footprint and then applying it to all transformer layers.
 When the sequence is too long, we cannot even profile one single transformer layer.
-In such extreme cases, we turn to the CUDA Unified Memory feature, which enables the swapping between GPU memory and CPU memory under the hook, effectively creating an illusion of unlimited GPU memory.
-By integrating CUDA Unified Memory support into the PyTorch CUDA allocator, we have successfully managed to profile the training of extremely long context lengths without encountering OOM errors.
-The profiler also gathers the basic information to determine $\alpha$ in Section 4.1, including the size of each skeletal activation tensor, and the forward time of a single transformer layer.
-Subsequently, it solves for the optimal $\alpha$ to maximize the overlapping of computation and tranmission as well as to avoid CPU OOM error.
+In such extreme cases, we turn to the CUDA Unified Memory feature, which enables the swapping between GPU memory and CPU memory under the hook, creating an illusion of unlimited GPU memory.
+By integrating CUDA Unified Memory support into the PyTorch CUDA allocator, we have successfully managed to profile the training of extremely long context lengths.
+The profiler also gathers the basic information to determine $\alpha$ in Section 4.1, including the size of each skeletal activation tensor, and the forward time of a one layer.
+Subsequently, it solves for the optimal $\alpha$ to maximize the overlapping of computation and transmission as well as to avoid CPU OOM error.
 4.3.3. Memory Planner
-Given the memory request sequence generated by the job profiler, memory planner executes the bi-level MIP optimization algorithm as introduced in Section 4.2 to generate a memory plan, which includes the address of each activation tensor and the peak memory usage needed during training.
+Given the memory request sequence generated by the job profiler, memory planner executes the bi-level MIP optimization algorithm as introduced in Section 4.2 to generate a memory plan, which includes the address of each transient activation tensor and the peak memory usage needed during training.
+Memo uses the Gurobi (Gurobi Optimization, LLC, 2024) optimizer to solve the MIP problems.
 In all our experiments, memory planning takes less than 5 minutes, which is negligible compared to the training time of LLMs.
@@
 4.3.4. Runtime Executor
-The runtime executor takes the memory plan as input, and executes the training process.
-It is built on the top of Megatron-LM (Shoeybi et al., 2019) and TransformerEngine (NVIDIA, 2024b), one of the most popular LLM training frameworks.
+The runtime executor takes the memory plan, and executes the training process.
+It is built on the top of Megatron-LM (Shoeybi et al., 2019) with TransformerEngine (NVIDIA, 2024b), one of the most popular training frameworks.
 The runtime executor utilizes two rounding buffers for the storage of skeletal activations, as introduced in Section 4.1.
-Meanwhile, the transient activation tensors are allocated and discarded according to the memory plan.
+Meanwhile, the transient activation tensors are (de)allocated according to the memory plan.
 Three CUDA streams are employed for efficient overlapping of data transmission and GPU computation, which are for GPU computation, activation offloading from GPU to CPU, and activation prefetching from CPU to GPU, respectively.
 Figure 10 shows the scheduling of computation and transmission.
-After the computation of a transformer layer’s forward pass, the skeletal activations of this layer are scheduled to be transferred to the CPU memory, which can overlap with the computation of the next layer.
-Before the backward computation of a transformer layer, the forward skeletal activations of the previous layer are scheduled to be fetched back to GPU.
-In addition, token-wise tensor recomputation is also scheduled before the layer’s backward pass.
-By hiding the activation swapping with computation and enabling the fractional, token-wise activation recomputation, Memo minimizes the overhead of activation rematerialization at full stretch.
-5. Experiments
-In this section, we conduct experiments across various model sizes and input sequence lengths to show that Memo achieves superior efficiency in longer context training of LLMs.
-5.1. Setup
-Hardware: Our experiments are conducted on an A800 GPU cluster, with each node equipped with 8 NVIDIA A800 GPUs (80GB).
-The GPUs within each node are interconnected via NVLinks (400GB/s), while the nodes are interconnected through Infiniband (200GB/s).
-Each node has 2TB CPU memory, and the GPU-CPU communication bandwidth is 32GB/s.
-Baselines: We select two widely-used LLM training frameworks as baselines for our experiments.
-The first is Megatron-LM (commit id: ccfeda47cb) (Shoeybi et al., 2019) in conjunction with TransformerEngine (v1.3) (NVIDIA, 2024b).
-Megatron-LM, maintained by NVIDIA, is renowned for its comprehensive support of hybrid parallelisms, including DP, TP, PP, SP, and CP.
-The other baseline is Megatron-Deepspeed (commit id 7eb36a11b3) paired with DeepSpeed (v0.14.3) (Rasley et al., 2020), which is recognized for ZeRO optimizers and DeepSpeed-Ulysses (Jacobs et al., 2023), a novel parallel training strategy designed for long context LLM training.
-Metrics: We use two important evaluation metrics to measure the training efficiency, which are Model FLOPs Utilization (MFU) and Tokens per GPU per Second (TGS).
-MFU is defined as the ratio of model FLOPs per second to the theoretical peak FLOPs per second of the GPU (e.g.
-312 TFLOPS for NVIDIA A800 GPUs) (Korthikanti et al., 2022).
-Based on FlashAttention (Dao, 2023) and considering the causal mask, the exact formula for calculating model FLOPS per sample is:
-$6\cdot s\cdot P+6\cdot n\cdot h\cdot s^{2}.$
-MFU is a standard metric that measures the training efficiency of how model FLOPs utilize computational resources.
-On the other hand, TGS directly measures training throughput, providing a clear view of how quickly a model can be trained using a given amount of training samples.
-Both metrics are crucial for LLM researchers and engineers, enabling comparisons among various training strategies (including distributed parallelisms and activation recomputation).
+After the computation of one layer’s forward pass, the skeletal activations of this layer are scheduled to be transferred to the CPU memory, which can overlap with the computation of the next layer.
+Before the backward computation of one layer, the forward skeletal activations of the previous layer are scheduled to be fetched back to GPU.
+In addition, token-wise tensor recomputation is scheduled before the layer’s backward pass.
+By hiding the activation swapping with computation and enabling the lightweight, token-wise activation recomputation, Memo minimizes the overhead of activation rematerialization at full stretch.
 Table 2. Configurations of the evaluated models.
@@
 65B 80 8192 32768 64 50257
+5. Experiments
+In this section, we conduct experiments across various model sizes and input sequence lengths to show that Memo achieves superior efficiency in long context training of LLMs.
+5.1. Setup
+Hardware: Our experiments are conducted on an A800 GPU cluster, with each node equipped with 8 NVIDIA A800 GPUs (80GB).
+The GPUs within each node are interconnected via NVLinks (400GB/s), while the nodes are interconnected through Infiniband (200GB/s).
+Each node has 2TB CPU memory, and the GPU-CPU communication bandwidth is 32GB/s.
+Baselines: We select two widely-used LLM training frameworks as baselines.
+The first is Megatron-LM (commit id: ccfeda47cb) (Shoeybi et al., 2019) in conjunction with TransformerEngine (v1.3) (NVIDIA, 2024b).
+Megatron-LM, maintained by NVIDIA, is renowned for its comprehensive support of hybrid parallelisms, including DP, TP, $P$, SP, and CP.
+The other baseline is Megatron-Deepspeed (commit id 7eb36a11b3) paired with DeepSpeed (v0.14.3) (Rasley et al., 2020), which is recognized for ZeRO optimizers and DeepSpeed-Ulysses (Jacobs et al., 2023), a novel parallel training strategy designed for long context LLM training.
+For all baselines, we employ FlashAttention (Dao et al., 2022; Dao, 2024) and mixed-precision training (Micikevicius et al., 2018; Kalamkar et al., 2019), which are commonly used in LLM training.
+Metrics: We use three important evaluation metrics to measure the training efficiency: Model FLOPs Utilization (MFU), Tokens per GPU per Second (TGS), and wall-clock time.
+MFU is defined as the ratio of model FLOPs per second to the theoretical peak FLOPs per second of the GPU (e.g.
+312 TFLOPS for NVIDIA A800 GPUs) (Korthikanti et al., 2022).
+Based on FlashAttention (Dao et al., 2022; Dao, 2024) and considering the causal mask, the exact formula for calculating model FLOPS per sample is:
+$6\cdot s\cdot P+6\cdot n\cdot h\cdot s^{2}.$
+MFU is a standard metric that measures the training efficiency of how model FLOPs utilize computational resources.
+On the other hand, TGS directly measures training throughput, and wall-clock time reflects the latency of training each batch, providing a clear view of how quickly a model can be trained using a given amount of training samples.
+All metrics are crucial for LLM researchers and engineers, enabling comparisons among various training strategies (including distributed parallelisms and activation recomputation).
 Workloads: Our experiments cover a wide range of workloads to examine the strength of Memo.
-In particular, we consider training the 7B, 13B, 30B and 65B GPT models on 8, 16, 32, and 64 GPUs, respectively, with various sequence lengths ranging from 64K to 1408K.
+In particular, we consider training the 7B, 13B, 30B and 65B GPT models on 8, 16, 32, and 64 GPUs, respectively, with various sequence lengths ranging from 4K to 1408K.
 The detailed model configurations are shown in Table 2.
 Table 3.
-MFU (Model FLOPS Utilization) and TGS (Throughput per GPU per Second) of different methods on different number of GPUs, model sizes, and sequence lengths.
+MFU (Model FLOPS Utilization), TGS (Tokens per GPU per Second), and wall-clock time of different methods on different numbers of GPUs, model sizes, and sequence lengths.
 “DS” stands for “DeepSpeed”, and “Mega” stands for “Megatron-LM”.
 ✗oom stands for Out Of Memory error when the system runs out of GPU memory, and ✗oohm stands for Out Of Host Memory error, which means the CPU memory is depleted.
-Sequence Length
-# GPUs Model Size Method 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
-27.95% 25.46% 23.38% oom oom oom oom oom oom oom oom oom
-DS 935.76 555.51 296.48 oom oom oom oom oom oom oom oom oom
-41.55% 24.13% 29.07% 27.98% 34.43% 30.90% oom oom oom oom oom oom
-8 7B Mega 1417.74 526.62 368.65 250.07 237.56 173.63 oom oom oom oom oom oom
-52.34% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71% 52.30% oom oom oom
-Memo 1786.22 1111.99 679.92 474.02 357.70 295.51 245.76 215.98 188.73 oom oom oom
-27.97% 25.45% 21.98% oom oom oom oom oom oom oom oom oom
-DS 553.85 333.51 171.82 oom oom oom oom oom oom oom oom oom
-38.51% 23.02% 25.30% 22.88% 29.10% 19.41% oom oom oom oom oom oom
-16 13B Mega 762.46 301.67 197.78 127.40 125.86 83.94 oom oom oom oom oom oom
-52.65% 50.93% 51.22% 51.91% 52.40% 52.13% 51.71% 51.76% 52.06% 51.74% 51.78% 52.10%
-Memo 1042.5 667.41 400.39 289.11 226.65 184.33 154.63 134.08 118.96 105.74 95.72 87.93
-29.93% 25.54% oom oom oom oom oom oom oom oom oom oom
-DS 296.41 176.92 oom oom oom oom oom oom oom oom oom oom
-35.76% 14.70% 17.15% 23.32% oom oom oom oom oom oom oom oom
-32 30B Mega 354.20 101.85 74.24 73.37 oom oom oom oom oom oom oom oom
-52.12% 49.66% 50.00% 50.69% 51.06% 51.72% 51.18% 51.50% 51.24% 51.73% 51.59% oohm
-Memo 516.16 344.09 216.41 159.52 126.23 105.28 88.55 77.48 68.20 61.72 55.79 oohm
-31.05% 26.13% 22.07% 20.40% 19.83% 19.06% 19.53% 19.12% 19.00% 19.11% 18.90% oom
-DS 149.8 90.15 48.51 32.89 25.24 20.04 17.50 14.93 13.14 11.86 10.64 oom
-22.79% 15.10% 9.57% 12.07% 5.32% oom oom oom oom oom oom oom
-64 65B Mega 109.94 52.10 21.03 19.46 6.77 oom oom oom oom oom oom oom
-47.80% 48.61% 49.87% 48.85% 49.71% 50.05% 51.16% 51.05% 51.27% 51.20% 51.42% 51.45%
-Memo 230.62 167.69 109.58 78.76 63.29 52.65 45.84 39.85 35.45 31.77 28.94 26.50
+“h”, “m” and “s” represent hour, minute and second respectively.
+7B GPT on 8 GPUs 13B GPT on 16 GPUs 30B GPT on 32 GPUs 65B GPT on 64 GPUs
+SeqLen DS Mega Memo DS Mega Memo DS Mega Memo DS Mega Memo
+44.71% 49.45% 49.45% 44.08% 46.17% 46.17% 43.70% 50.67% 50.67% 43.89% 53.61% 53.61%
+3235.64 3578.86 3578.86 1675.53 1755.23 1755.23 724.44 840.14 840.14 337.95 412.90 412.90
+4K 2.53s 2.29s 2.29s 2.44s 2.33s 2.33s 2.83s 2.44s 2.44s 3.03s 2.48s 2.48s
+46.28% 51.57% 51.57% 38.90% 50.91% 50.91% 38.78% 49.27% 49.27% 43.34% 53.27% 53.27%
+3116.25 3472.95 3472.95 1393.24 1823.52 1823.52 615.22 781.68 781.68 321.00 394.61 394.61
+8K 5.26s 4.72s 4.72s 5.88s 4.50s 4.50s 6.66s 5.24s 5.24s 6.38s 5.19s 5.19s
+36.41% 51.29% 51.29% 36.95% 51.52% 51.52% 34.54% 49.84% 49.84% 40.98% 51.10% 51.10%
+2152.65 3032.14 3032.14 1186.24 1654.11 1654.11 504.64 728.22 728.22 282.05 351.58 351.58
+16K 15.2s 10.8s 10.8s 13.8s 9.91s 9.91s 16.2s 11.2s 11.2s 14.5s 11.6s 11.6s
+35.91% 52.75% 52.75% 34.51% 49.79% 49.79% 35.19% 48.06% 48.06% 35.73% 34.00% 44.79%
+1706.13 2506.39 2506.39 917.96 1324.27 1324.27 443.77 606.16 606.16 215.32 204.83 269.84
+32K 38.4s 26.1s 26.1s 35.7s 24.7s 24.7s 36.9s 27.0s 27.0s 38.0s 39.9s 30.4s
+27.95% 41.55% 52.34% 27.97% 38.51% 52.65% 29.93% 35.76% 52.12% 31.05% 22.79% 47.80%
+935.76 1417.74 1786.22 553.85 762.46 1042.50 296.41 354.20 516.16 149.80 109.94 230.62
+64K 2m17s 1m33s 1m13s 1m58s 1m26s 1m3s 1m51s 1m33s 1m3s 1m49s 2m19s 1m11s
+25.46% 24.13% 50.96% 25.45% 23.02% 50.93% 25.54% 14.70% 49.66% 26.13% 15.10% 48.61%
+555.51 526.62 1111.99 333.51 301.67 667.41 176.92 101.85 344.09 90.15 52.10 167.69
+128K 7m52s 8m18s 3m56s 6m33s 7m15s 3m16s 6m10s 10m43s 3m10s 6m4s 10m30s 3m15s
+23.38% 29.07% 53.62% 21.98% 25.30% 51.22% oom 17.15% 50.00% 22.07% 9.57% 49.87%
+296.48 368.65 679.92 171.82 197.78 400.39 oom 72.24 216.41 48.51 21.03 109.58
+256K 29m28s 23m42s 12m51s 25m26s 22m5s 10m55s oom 29m30s 10m6s 22m31s 51m56s 9m58s
+oom 27.98% 53.04% oom 22.88% 51.91% oom 23.32% 50.69% 20.40% 12.07% 48.85%
+oom 250.07 474.02 oom 127.40 289.11 oom 73.37 159.52 32.89 19.46 78.76
+384K oom 52m25s 27m39s oom 51m26s 22m40s oom 44m39s 20m32s 49m49s 1h24m 20m48s
+oom 34.43% 51.84% oom 29.10% 52.40% oom oom 51.06% 19.83% 5.32% 49.71%
+oom 237.56 357.70 oom 125.86 226.65 oom oom 126.23 25.24 6.77 63.29
+512K oom 1h14m 48m51s oom 1h10m 38m33s oom oom 34m37s 1h27m 5h23m 34m31s
+oom 30.90% 52.59% oom 19.41% 52.13% oom oom 51.72% 19.06% oom 50.05%
+oom 173.63 295.51 oom 83.94 184.33 oom oom 105.28 20.04 oom 52.65
+640K oom 2h6m 1h14m oom 2h10s 59m15s oom oom 51m52s 2h16m oom 51m52s
+oom oom 51.89% oom oom 51.71% oom oom 51.18% 19.53% oom 51.16%
+oom oom 245.76 oom oom 154.63 oom oom 88.55 17.50 oom 45.84
+768K oom oom 1h47m oom oom 1h25m oom oom 1h14m 3h7m oom 1h11m
+oom oom 52.71% oom oom 51.76% oom oom 51.50% 19.12% oom 51.05%
+oom oom 251.98 oom oom 134.08 oom oom 77.48 14.93 oom 39.85
+896K oom oom 2h22m oom oom 1h54m oom oom 1h39m 4h16m oom 1h36m
+oom oom 52.30% oom oom 52.06% oom oom 51.24% 19.00% oom 51.27%
+oom oom 188.73 oom oom 118.96 oom oom 68.20 13.14 oom 35.45
+1024K oom oom 3h5m oom oom 2h27m oom oom 2h8m 5h33m oom 2h3m
+oom oom oom oom oom 51.74% oom oom 51.73% 19.11% oom 51.20%
+oom oom oom oom oom 105.74 oom oom 61.72 11.86 oom 31.77
+1152K oom oom oom oom oom 3h6m oom oom 2h39m 6h43m oom 2h35m
+oom oom oom oom oom 51.78% oom oom 51.59% 18.90% oom 51.42%
+oom oom oom oom oom 95.72 oom oom 55.79 10.64 oom 28.94
+1280K oom oom oom oom oom 3h48m oom oom 3h16m 8h33m oom 3h9m
+oom oom oom oom oom 52.10% oom oom oohm oom oom 51.45%
+oom oom oom oom oom 87.93 oom oom oohm oom oom 26.50
+1408K oom oom oom oom oom 4h33m oom oom oohm oom oom 3h47m
 5.2. End-to-end Evaluation
 We compare the end-to-end training efficiency of Memo and two baselines.
-Table 3 shows the MFU and TGS of DeepSpeed-Ulysses, Megatron-LM and Memo under different training workloads.
+Table 3 shows the MFU, TGS, and wall-clock time of DeepSpeed-Ulysses, Megatron-LM and Memo under different training workloads.
 During evaluation, we manually adjust the distributed parallelism strategies for each system and each workload to achieve optimal training performance for fair comparisons.
-Detailed parallelism strategies are provided in Appendix A.
 Overall, Memo is capable of training longer sequences than the competitors.
@@
 When training smaller models, DeepSpeed supports only very small sequence lengths.
-This is because it can only utilize a small SP size of 8, which either aligns to the number of GPUs or is dividable by the number of attention heads (40 and 56).
+This is because it can only utilize a small SP degree of 8, which either aligns to the number of GPUs or is dividable by the number of attention heads (40 and 56).
 In contrast, by token-wise recomputation/swapping and memory planning, Memo is able to train sequences over 1 million tokens in all scenarios.
-Furthermore, when comparing Memo to the baselines with aligned sequence lengths, Memo achieves superior MFU and TGS.
-Across all experimented workloads, Memo achieves an average MFU of 51.33%.
-In contrast, Megatron-LM and DeepSpeed only achieve an average MFU of 23.91% and 23.26%, respectively.
-In average, Memo achieves $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
-The deficiencies of baselines are not surprising — due to their unsatisfactory memory management, Megatron-LM and DeepSpeed require high model parallel size (i.e.
-large TP size and/or large SP size) and full activation recomputation to avoid OOM error, which leads to low training efficiency.
-For example, in order to train the 65B model with 1024K sequence length on 64 GPUs, DeepSpeed has to simultaneously use an SP size of 64, ZeRO-3 and activation recomputation to avoid OOM error.
-The large SP/TP size and ZeRO degree introduce significant communication overheads, while vanilla memory reduction techniques incurs extra overhead, further diminishing training efficiency.
-Memory fragmentation is another factor leading to the low training efficiency of these systems.
-When the GPU memory becomes highly fragmented and cannot allocate sufficient space for new tensors, the allocator will call expensive “cudaMalloc” and “cudaFree” operations to reorganize the GPU memory, which blocks GPU computation substantially.
+When training with relatively short sequences ranging from 4K to 32K, there’s little memory pressure.
+In such cases, activation recomputation is often unnecessary, and different methods may use the same parallelism strategies.
+Consequently, Memo falls back to Megatron-LM, resulting in the same performance.
+Nevertheless, Memo significantly outperforms the baselines when the sequence length reaches 64K or even higher.
+Furthermore, when comparing Memo to the baselines with aligned sequence lengths, Memo achieves superior MFU, TGS, and wall-clock time.
+Across all experimented workloads, Memo achieves an average MFU of 51.04%.
+In contrast, Megatron-LM and DeepSpeed only achieve an average MFU of 35.01% and 30.74%, respectively.
+On average, Memo achieves $1.97\times$ and $1.80\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
+The deficiencies of baselines are not surprising — due to the unsatisfactory memory management, they are forced to utilize less-efficient configurations of parallelism strategies and rematerialization options.
+[^6] We have provided the detailed configurations in our supplementary materials.
+Firstly, the baselines need to employ a high model parallelism degree (i.e.
+large TP degree and/or large SP degree) to avoid OOM error.
+For instance, when training the 65B model with 256K sequence length, Megatron-LM has to use a TP degree of 16, while Memo can support a TP degree of 8.
+Given the much lower inter-node bandwidth compared to intra-node, the TP communication cost of Memo is 63% lower than that of Megatron-LM (44.3 v.s.
+120.4 seconds).
+Secondly, the baselines require full activation recomputation to avoid the OOM error, leading to over 20% redundant GPU calculation.
+In contrast, Memo integrates fine-grained recomputation and swapping, incurring smaller overhead.
+Last but not least, the lack of memory planning often triggers memory reorganization during training, which blocks GPU computation and significantly hampers training efficiency.
 When training the 7B model on 8 GPUs using Megatron-LM, the memory reorganization operation is triggered 6 times and 16 times per iteration for sequence lengths of 128K and 256K, respectively.
-In contrast, Memo successfully addresses these issues.
-On the one hand, the fine-grained activation recomputation and swapping technique significantly reduces computation burden compared to the vanilla full recomputation.
-This advantage is even more pronounced during long-context training.
-On the other hand, Memo adopts an optimized memory planning for activation, which avoids GPU memory fragmentation and the time-consuming memory reorganization process.
-Reducing memory fragmentation also allows for more efficient model parallelism configurations.
-For example, when training the 30B model on 32 GPUs with 256k sequence length, Memo adopts a TP size of 8, a CP size of 2, and a DP size of 2, which is more efficient than Megatron-LM with a TP size of 8 and CP size of 4.
-Additionally, Memo does not trigger any memory reorganization since the memory is already managed.
-As a result, Memo consistently achieves an MFU of approximately 50% across all model sizes and sequence lengths, enabling the efficient training of significantly longer sequences compared to the baselines.
+A detailed ablation study is presented in Section 5.3.
+As a result, Memo consistently achieves an MFU of approximately 50% across all model sizes and sequence lengths, enabling more efficient training of significantly long sequences compared to the baselines.
 Table 4. MFU of different methods for ablation studies. The experiments are conducted by training the 7B model on 8 GPUs.
@@
 Full Swapping + Memory Plan 37.40% 46.33% 53.62% oohm oohm oohm oohm oohm
-Memo (Fine-grained Recomputation & Swapping + Memory Plan) 47.99% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71%
+Memo (Fine-grained Management + Memory Plan) 47.99% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71%
+Table 5. Impact of offloading fraction $\alpha$ on training efficiency.
+Value of $\alpha$
+Sequence Length 0.000 0.125 0.250 0.375 0.500 0.625 0.750 0.875 1.000
+192K 51.04% 51.86% 52.14% 52.30% 52.72% 52.78% 53.11% 51.74% 49.69%
+256K 52.03% 52.05% 52.19% 52.45% 52.70% 52.96% 53.21% 53.29% 53.62%
+320K 52.15% 52.36% 52.52% 52.66% 52.92% 53.10% 53.35% oohm oohm
+384K 52.09% 52.31% 52.45% 52.58% 53.04% oohm oohm oohm oohm
+5.3. Ablation Studies
+Next, we assess the effectiveness of the proposed techniques in Memo.
+All experiments in ablation studies are conducted by training the 7B model on 8 GPUs, keeping the parallelism configuration fixed at a TP degree of 4 and a CP degree of 2.
+5.3.1. Effectiveness of Memory Planning
+To evaluate the effectiveness of memory planning, we evaluate two variants of Memo with full recomputation, both with and without memory planning.
+As shown in the first two rows of Table 4, without memory planning, the longest sequence supported is only 384K, achieving an MFU of 25.67%.
+After applying memory planning, the longest supported sequence length increases to 640K, with an MFU of 42.15%.
+The results are reasonable since full recomputation without memory planning has severe memory fragmentation, resulting in OOM errors in large sequence length scenarios.
+Additionally, the frequent GPU memory reorganization process further impairs training efficiency.
+By employing memory planning, the fragmentation issue can be eliminated, providing more memory for longer context training.
+Getting rid of GPU memory reorganization, memory planning brings an average of $1.51\times$ MFU when facing the same context length.
+5.3.2. Effectiveness of Token-wise Recomputation
+For token-wise recomputation and swapping, we compare Memo and its variants, one with full recomputation and another with full swapping.
+The results are shown in the last three rows of Table 4.
+When training with appropriate sequence length, which is 256k in this scenario, the computation time of one transformer layer can fully overlap with the offloading time of a layer’s activations.
+Therefore, full swapping with memory planning can achieve an MFU of 53.62% under 256K sequence length, far exceeding the 42.05% achieved by full recomputation with memory planning.
+However, for short sequence lengths, such as 64K, the offloading time of one layer’s activations blocks the GPU computation, resulting in a lower MFU than full recomputation.
+Full swapping presents another challenge as the sequence length grows longer: the host memory is rapidly depleted by offloaded activations, leading to OOHM errors.
+Full recomputation, which does not offload the input tensor of each layer, also encounters OOM errors at a sequence length of 768K.
+By employing token-wise recomputation together with swapping, Memo consistently improves training efficiency for both short and long context lengths.
+For short sequence lengths like 64K, our tensor-level design only offloads the input tensor of the transformer layer and the FlashAttention output tensor to CPU memory, enabling efficient overlap of GPU computation and data transmission.
+For long context lengths, our rounding-buffer token-level management successfully avoids depleting the CPU memory, and incurs only minimal recomputation overhead.
+Among all the methods, Memo supports the longest sequence length.
+Considering MFU, Memo achieves an average of $1.22\times$ MFU compared to full recomptutation with memory planning, and an average of $1.13\times$ MFU compared to full offloading with memory planning.
+5.3.3. Effect of the Offloading Fraction $\alpha$
+To investigate the impact of the offloading fraction $\alpha$ on training efficiency, we train a 7B model on 8 GPUs, with TP=4=4, CP=2=2 and varying $\alpha$ values.
+The results are presented in Table 5.
+For a sequence length of 192K, the MFU initially increases with $\alpha$, peaking at 53.11% when $\alpha=0.75$, and then declines.
+This trend occurs because, for $\alpha<0.75$, a higher $\alpha$ reduces recomputation overhead by increasing the proportion of skeletal activations that are swapped.
+Beyond this point, however, further increases in $\alpha$ lead to swapping overhead that exceeds the computation time of an individual transformer layer, stalling GPU computations until the swapping concludes and thus reducing the MFU.
+At a sequence length of 256K, the computation time of a single transformer layer, due to the quadratic complexity of self-attention, surpasses the offloading time for all skeletal activations in this layer.
+Consequently, training efficiency consistently improves as $\alpha$ increases.
+For longer sequences such as 320K and 384K, offloading all skeletal activations becomes unviable due to CPU memory constraints.
+Considering the constraints, Memo can effectively identify the optimal $\alpha$ to minimize recomputation overhead.
 (a) Longest supported sequence length.
@@
 (d): Training loss curves of Megatron-LM and Memo with different $\alpha$.
-5.3. Ablation Studies
-Next, we assess the effectiveness of the proposed techniques in Memo.
-All experiments in ablation studies are conducted by training the 7B model on 8 GPUs, keeping the parallelism configuration fixed at a TP size of 4 and a CP size of 2.
-5.3.1. Effectiveness of Memory Planning
-To evaluate the effectiveness of memory planning, we evaluate two variants of Memo with full recomputation, both with and without memory planning.
-As shown in the first two rows of Table 4, without memory planning, the longest sequence supported is only 384k, achieving an MFU of 25.67%.
-After applying memory planning, the longest supported sequence length increases to 640k, with an MFU of 42.15%.
-The results are reasonable since full recomputation without memory planning has severe memory fragmentation, resulting in OOM errors in large sequence length scenarios.
-Additionally, the frequent GPU memory reorganization process further impairs training efficiency.
-By employing memory planning, the fragmentation issue can be minimized, providing more memory for longer context training.
-Getting rid of GPU memory reorganization, memory planning brings an average of $1.51\times$ MFU when facing the same context length.
-5.3.2. Effectiveness of Token-wise Recomputation
-For token-wise recomputation and swapping, we compare Memo and its variants, one with full recomputation and another with full swapping.
-The results are shown in the last two rows of Table 4.
-When training with appropriate sequence length, which is 256k in this scenario, the computation time of one transformer layer can fully overlap with the offloading time of a layer’s activations.
-Therefore, full swapping with memory planning can achieve an MFU of 53.62% under 256K sequence length, far exceeding the 42.05% achieved by full recomputation with memory planning.
-However, for short sequence lengths, such as 64K, the offloading time of one layer’s activations block the GPU computation, resulting in a lower MFU than full recomputation.
-Full swapping presents another challenge as the sequence length grows longer: the host memory is rapidly depleted by offloaded activations, leading to OOHM errors.
-By employing token-wise recomputation together with swapping, Memo consistently improves training efficiency for both short and long context lengths.
-For short sequence lengths like 64K, our tensor-level design only offloads the input tensor of the transformer layer and the FlashAttention output tensor to CPU memory, enabling efficient overlap of GPU computation and data transmission.
-For long context lengths, our token-level management successfully avoids depleting the CPU memory, and incurs only minimal recomputation overhead.
-Among all the methods, Memo supports the longest sequence length.
-Considering MFU, Memo achieves an average of $1.22\times$ MFU compared to full recomptutation with memory planning, and an average of $1.13\times$ MFU compared to full offloading with memory planning.
 5.4. Scalability
-To demonstrate the scalability of Memo, we train the 7B model on 8, 16, 32, and 64 GPUs respectively, and report the maximum supported sequence length.
+To show the scalability of Memo, we train the 7B model on 8, 16, 32, and 64 GPUs respectively, and report the maximum supported sequence length.
 As shown in Figure 11(a), when the number of GPUs increases, the maximum sequence length supported by Memo increases linearly.
@@
 Memo also consistently maintains an MFU of over 50% across different numbers of GPUs, as shown in Figure 11(b).
-For DeepSpeed, as the number of GPUs increases, it can enlarge the SP size, leading to longer supported sequence lengths.
-Note that since because the 7B model has 32 attention heads and thereby the maximum SP size is 32, DeepSpeed achieves the same maximum sequence length 1536K for both 32 and 64 GPUs.
-Megatron-LM supports context parallelism, which has better scalability than DeepSpeed.
+For DeepSpeed, as the number of GPUs increases, it can enlarge the SP degree, leading to longer supported sequence length.
+Note that since the 7B model has 32 attention heads, the maximum SP degree is 32.
+As a result, DeepSpeed achieves the same maximum sequence length of 1536K for both 32 and 64 GPUs.
+Megatron-LM supports context parallelism, which has better scalability.
 When the number of GPUs increases, the longest sequence length it can handle grows sublinearly.
 Compared to the baselines, Memo introduces fine-grained activation memory management, achieving not only ideal scalability but also better efficiency.
-We also evaluate the MFU metrics of the three systems when training the 7B model on 64 GPUs with sequence lengths varing from 1024K to 8192K.
-In Figure 12(c), as sequence length increases, the MFU of Memo maintains above 50% consistently, surpassing the competitors significantly.
-5.5. Convergence of Memo
-To demonstrate the correctness of our system implementation, we conduct a convergence experiment.
-Specifically, we train the 7B model with 128K sequence length on 8 GPUs and compare the convergence of Memo and Megatron-LM for 1000 iterations.
-For both systems, we fix the parallelism strategy to TP size 4 and CP size 2, and for Memo, we enumerate the value of $\alpha$ in $\{0,0.125,0.25,0.5,1\}$.
-As shown in Figure 12(d), the loss curves of Memo with different $\alpha$ values all align with Megatron-LM, confirming the correctness of our system implementation.
+We also evaluate the MFU of the three systems when training the 7B model on 64 GPUs with sequence lengths varing from 1024K to 8192K.
+In Figure 11(c), as sequence length increases, the MFU of Memo maintains above 50%, surpassing the baselines.
+$5.$ Convergence of Memo
+To confirm the correctness of our system implementation, we conduct a convergence experiment.
+We train a 7B model with 128K sequence length on 8 GPUs and compare the convergence of Memo and Megatron-LM for 1000 iterations.
+For both systems, we fix the parallelism strategy to TP=4=4 and CP=2=2.
+For Memo, we enumerate the value of $\alpha$ in $\{0,0.125,0.25,0.5,1\}$.
+As shown in Figure 11(d), the loss curves of Memo with different $\alpha$ values all align with Megatron-LM, confirming the correctness of Memo.
 6. Related Work
+Data management and machine learning: Recently, the intersection of data management and machine learning has gained popularity, with numerous studies emerging from the data management community (Miao et al., 2024a, 2023; Zhang et al., 2022; Zhao et al., 2023; Murray et al., 2021; Um et al., 2023; Isenko et al., 2022; Mohan et al., 2021; Miao et al., 2021; Li et al., 2023b; Gao et al., 2024; Miao et al., 2022; Wang et al., 2024; Yuan et al., 2021; Khamis et al., 2020; Nakandala et al., 2020; Koutsoukos et al., 2021; Schleich et al., 2023; Zhao et al., 2022; Zhou et al., 2024; Wang et al., 2023; Yu et al., 2023; Ooi et al., 2024; Zhang et al., 2023b, 2024a).
+Some works focus on optimizing I/O data management to accelerate data pre-processing pipelines (Murray et al., 2021; Um et al., 2023; Isenko et al., 2022; Mohan et al., 2021).
+Another work proposes to use a fine-grained automatic parallelism strategy to speed up model training (Miao et al., 2022).
+Memo also employs data management techniques and methodologies to enhance the efficiency of machine learning tasks, specifically addressing the challenges associated with long-sequence LLM training.
 Parallelism strategies for long context training: To tackle the challenge of long context training, DeepSpeed-Ulysses (Jacobs et al., 2023) employs an novel AllToAll communication to facilitate the partition of input sequence among GPUs, achieving lower communication overhead compared with Megatron-LM sequence parallelism (Korthikanti et al., 2022).
-LightSeq (Li et al., 2023), Ring Attention (Liu et al., 2023b), and Megatron-LM context parallelism (NVIDIA, 2024a) propose to split the sequence within self-attention computation, achieving better scalability.
-Recent efforts in the realm of sequence and context parallelisms (Chen et al., 2024; Fang and Zhao, 2024) aim to integrate multiple strategies and enhance existing distributed settings.
+LightSeq (Li et al., 2023a), Ring Attention (Liu et al., 2023), and Megatron-LM context parallelism (NVIDIA, 2024a) propose to split the sequence within self-attention computation, achieving better scalability.
+Recent efforts in the realm of sequence and context parallelisms (Chen et al., 2024; Fang and Zhao, 2024; Ge et al., 2024) aim to integrate multiple strategies and enhance existing distributed settings.
 It is worth noting that the fine-grained memory management of Memo is orthogonal to these distributed parallelism strategies.
@@
 In addition to tensor recomputation, MegTaichi (Hu et al., 2022) also proposes to co-optimize the tensor partition.
-Coop (Zhang et al., 2023) notices that naive tensor recomputation leads to severe memory fragmentation, and proposes heuristics to reduce memory fragmentation during tensor recomputation.
+Coop (Zhang et al., 2023a) notices that naive tensor recomputation leads to severe memory fragmentation, and proposes heuristics to reduce memory fragmentation during tensor recomputation.
 While these works offer solutions for common deep learning models, they do not take advantage of the specific characteristics of LLM training to achieve full overlapping and fragmentation minimization.
-Memory planning for deep learning models: The memory allocation problem of in deep learning models can be regarded as a DSA problem and solved by MIP (Sekiyama et al., 2018).OLLA (Steiner et al., 2022) proposes to optimize the lifetime and memory location of tensors during the training process by solving a joint ILP problem, reducing the peak memory during training.
-However, these methods does not exploit the repetitive substructure in LLMs and relies on heuristics to simplify the integer programming problem.
+Memory planning for deep learning models: The memory allocation problem in deep learning models can be regarded as a DSA problem and solved by MIP (Sekiyama et al., 2018).
+OLLA (Steiner et al., 2022) proposes to optimize the lifetime and memory location of tensors during the training process by solving a joint ILP problem, reducing the peak memory during training.
+However, it does not exploit the repetitive substructure in LLMs and relies on heuristics to simplify the integer programming problem.
 7. Conclusion
@@
 We designed a fine-grained activation recomputation and swapping strategy to fully utilize the idle PCIe bandwidth during the GPU computation, thereby reducing the activation rematerialization cost in long context LLM training.
-We employed a bi-level MIP technique to solve the problem of memory allocation within one transformer layer, and reused the same memory space for each identical layer so as to eliminate memory fragmentation.
-Through extensive experiments, we demonstrated that Memo achieved an average of $2.42\times$ MFU compared to Megatron-LM.
+We employed a bi-level MIP technique to first solve the problem of memory allocation within one transformer layer, and then reused the same memory space for each identical layer so as to eliminate memory fragmentation.
+Through extensive experiments, we demonstrated that Memo achieved an average of $1.97\times$ MFU compared to Megatron-LM.
 By leveraging fine-grained tensor memory management, Memo achieved 52.30% MFU when training 7B LLM with 1 million sequence length on only 8 A800 GPUs.
-Appendix A Detailed Parallelism Strategy in Evaluation
-Table 56, and 7 present the detailed parallelism training strategies for end-to-end evaluation.
-Table 5.
-Detailed parallelism strategies for DeepSpeed.
-“SP” stands for DeepSpeed-Ulysses sequence parallel size, “DP” stands for data parallel size, “ZeRO” stands for ZeRO stage, “AR” stands for activation recomputation.
-Sequence Length
-# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
-SP 2 4 8 oom oom oom oom oom oom oom oom oom
-DP 4 2 1 oom oom oom oom oom oom oom oom oom
-ZeRO 3 3 3 oom oom oom oom oom oom oom oom oom
-8 7B AR On On On oom oom oom oom oom oom oom oom oom
-SP 4 8 8 oom oom oom oom oom oom oom oom oom
-DP 4 2 2 oom oom oom oom oom oom oom oom oom
-ZeRO 3 3 3 oom oom oom oom oom oom oom oom oom
-16 13B AR On On On oom oom oom oom oom oom oom oom oom
-SP 8 8 oom oom oom oom oom oom oom oom oom oom
-DP 4 4 oom oom oom oom oom oom oom oom oom oom
-ZeRO 3 3 oom oom oom oom oom oom oom oom oom oom
-32 30B AR On On oom oom oom oom oom oom oom oom oom oom
-SP 8 8 16 32 32 32 64 64 64 64 64 oom
-DP 8 8 4 2 2 2 1 1 1 1 1 oom
-ZeRO 3 3 3 3 3 3 3 3 3 3 3 oom
-64 65B AR On On On On On On On On On On On oom
-Table 6.
-Detailed parallelism strategies for Megatron-LM.
-“TP” stands for tensor parallel size, “CP” stands for context parallel size, “DP” stands for data parallel size, “PP” stands for pipeline parallel size, “AR” stands for activation recomputation.
-For all evaluations of Megatron-LM, Megatron-style sequence parallel and ZeRO-1 optimizer are enabled.
-Sequence Length
-# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
-TP 2 4 4 4 oom oom oom oom oom oom oom oom
-CP 4 2 2 2 oom oom oom oom oom oom oom oom
-DP 1 1 1 1 oom oom oom oom oom oom oom oom
-$P$ 1 1 1 1 oom oom oom oom oom oom oom oom
-8 7B AR On On On On oom oom oom oom oom oom oom oom
-TP 8 4 4 4 8 8 oom oom oom oom oom oom
-CP 2 2 2 2 2 2 oom oom oom oom oom oom
-DP 1 1 1 1 1 1 oom oom oom oom oom oom
-$P$ 1 2 2 2 1 1 oom oom oom oom oom oom
-16 13B AR On On On On On On oom oom oom oom oom oom
-TP 4 4 8 8 oom oom oom oom oom oom oom oom
-CP 4 8 4 4 oom oom oom oom oom oom oom oom
-DP 1 1 1 1 oom oom oom oom oom oom oom oom
-$P$ 2 1 1 1 oom oom oom oom oom oom oom oom
-32 30B AR On On On On oom oom oom oom oom oom oom oom
-TP 8 32 16 16 64 oom oom oom oom oom oom oom
-CP 8 1 1 4 1 oom oom oom oom oom oom oom
-DP 1 2 1 1 1 oom oom oom oom oom oom oom
-$P$ 1 1 4 1 1 oom oom oom oom oom oom oom
-64 65B AR On On On On On oom oom oom oom oom oom oom
-Table 7.
-Detailed parallelism strategies for Memo.
-“TP” stands for tensor parallel size, “CP” stands for context parallel size, “DP” stands for data parallel size, “PP” stands for pipeline parallel size, $\alpha$ stands for CPU offload fraction.
-For all evaluations of Memo, Megatron-style sequence parallel and ZeRO-1 optimizer are enabled.
-Sequence Length
-# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
-TP 2 4 4 4 4 4 4 4 8 oom oom oom
-CP 2 2 2 2 2 2 2 2 1 oom oom oom
-DP 2 1 1 1 1 1 1 1 1 oom oom oom
-$P$ 1 1 1 1 1 1 1 1 1 oom oom oom
-8 7B $\alpha$ 0.0 0.0 1.0 0.5 0.125 0.125 0.0 0.0 0.0 oom oom oom
-TP 2 4 4 4 4 4 4 4 4 4 4 4
-CP 2 2 2 2 4 4 4 4 4 4 4 4
-DP 4 2 2 1 1 1 1 1 1 1 1 1
-$P$ 1 1 1 2 1 1 1 1 1 1 1 1
-16 13B $\alpha$ 0.0 0.0 0.0 0.0 0.5 0.125 0.125 0.0 0.0 0.0 0.0 0.0
-TP 4 8 8 8 8 8 8 8 8 8 8 oohm
-CP 1 1 2 2 2 4 4 4 4 4 4 oohm
-DP 8 4 2 2 2 1 1 1 1 1 1 oohm
-$P$ 1 1 1 1 1 1 1 1 1 1 1 oohm
-32 30B $\alpha$ 0.0 0.0 0.0 0.0 0.0 0.25 0.125 0.125 0.125 0.125 0.0 oohm
-TP 8 4 8 8 8 8 8 8 8 8 8 8
-CP 1 4 4 8 8 8 8 8 8 8 8 8
-DP 8 4 2 1 1 1 1 1 1 1 1 1
-$P$ 1 1 1 1 1 1 1 1 1 1 1 1
-64 65B AR 0.0 0.0 0.0 0.25 0.25 0.25 0.25 0.125 0.125 0.0 0.0 0.0
+Acknowledgement
+This work is supported by National Science and Technology Major Project (2022ZD0116315), National Natural Science Foundation of China (U22B2037, U23B2048, 62402011), Beijing Municipal Science and Technology Project (Z231100010323002), China National Postdoctoral Program for Innovative Talents (BX20230012), China Postdoctoral Science Foundation (2024M750103), Beijing Natural Science Foundation (4244080), research grant No.
+IPT-2024JK29, PKU-Tencent joint research Lab, and High-performance Computing Platform of Peking University.
+Fangcheng Fu and Bin Cui are the corresponding authors.